### v1

# OpeNLGauge: An Explainable Metric for NLG Evaluation with Open-Weights LLMs

Abstract
Large Language Models (LLMs) have demonstrated great potential as evaluators of NLG systems, allowing for high-quality, reference-free, and multi-aspect assessments. However, existing LLM-based metrics suffer from two major drawbacks: reliance on proprietary models to generate training data or perform evaluations, and a lack of fine-grained, explanatory feedback. In this paper, we introduce OpeNLGauge, a fully open-source, reference-free NLG evaluation metric that provides accurate explanations based on error spans. OpeNLGauge is available as a two-stage ensemble of larger open-weight LLMs, or as a small fine-tuned evaluation model, with confirmed generalizability to unseen tasks, domains and aspects. Our extensive meta-evaluation shows that OpeNLGauge achieves competitive correlation with human judgments, outperforming state-of-the-art models on certain tasks while maintaining full reproducibility and providing explanations more than twice as accurate.
OpeNLGauge: An Explainable Metric for NLG Evaluation with Open-Weights LLMs
Ivan Kartáč  and  Mateusz Lango  and  Ondřej Dušek Charles University, Faculty of Mathematics and Physics Institute of Formal and Applied Linguistics Prague, Czech Republic ivan.kartac@gmail.com, {lango,odusek}@ufal.mff.cuni.cz
1 Introduction
Evaluating Natural Language Generation (NLG) systems remains a challenging research problem. Traditional overlap-based metrics, such as BLEU Papineni et al. (2002a) and ROUGE Lin (2004a), are still widely used but exhibit limited correlation with human judgments, particularly when assessing modern NLG systems Novikova et al. (2017a). With the rise of pre-trained language models, the research community began to shift toward model-based metrics that better capture semantic similarity, yet their performance was still unsatisfactory Yan et al. (2023); Glushkova et al. (2023).
Figure 1: Overview of the proposed approach: OpeNLGaugeens ensemble and its distilled version OpeNLGaugeft.
Recently, Large Language Models (LLMs) have demonstrated remarkable potential in imitating human evaluation of generated text Jiang et al. (2023); Xu et al. (2023a); Hu et al. (2024b). LLM-based metrics are often general enough to evaluate diverse NLG tasks and can provide high evaluation performance without the need for reference texts Liu et al. (2023). They also offer better adaptability to evaluate specific aspects of generated text and are able to evaluate beyond semantic correctness, taking into account aspects like factual consistency or relevance Gu et al. (2025). However, existing LLM-based NLG metrics lack two important features: (1) they are not fully open-source, with most relying on data generated by proprietary LLMs such as GPT-4 OpenAI et al. (2024), and (2) they do not provide precise explanations for their evaluations, with most of them producing only overall scores, or explanations limited to a single short comment (see Table 1 for details).
In this paper, we introduceOpeNLGauge – a versatile, reference-free metric for NLG tasks that provides precise, error-span-based explanations (Figure 1). Unlike existing LLM-based metrics, OpeNLGauge is built entirely on open-weight models and does not rely on annotated datasets. It supports a wide range of NLG tasks, including data-to-text, summarization or story generation. Moreover, it allows for fine-grained evaluation of a specific, customizable aspect of the generated text like faithfulness, fluency or coherence. Our contributions are as follows:
  • We propose an effective prompting strategy and a two-stage LLM ensemble to obtain high-quality evaluations of NLG outputs. Interestingly, the proposed OpeNLGaugeens demonstrates a higher correlation with human judgements on some NLG tasks than previously proposed metrics based on proprietary LLMs.
  • We collect outputs from over a hundred different NLG systems on 15 NLG tasks and use OpeNLGaugeens to annotate them to construct a comprehensive dataset for training open NLG metrics.
  • We use the collected dataset to construct OpeNLGaugeft – a fine-tuned version of the smallest model from the Llama 3.1 family Grattafiori et al. (2024), which is able to provide accurate explanations and good quality assessments of NLG outputs in a more cost-effective way.
  • We perform an extensive meta-evaluation of the proposed metrics on seven datasets covering five NLG tasks. The experimental analysis includes human evaluation of explanation quality, several ablation experiments, and evaluation on tasks, domains and evaluation aspects not seen during training.
Our experiments show that OpeNLGauge achieves higher explanation quality than other metrics trained on data generated by proprietary LLMs, while also delivering strong evaluation performance. On tasks such as summarization, OpeNLGauge achieves higher correlations with human judgments than the strong proprietary GPT-4 model. While the state-of-the-art metric, Themis – which leverages both human-annotated data and data from proprietary LLMs – remains superior overall, OpeNLGauge proves highly competitive and even outperforms it on certain tasks, while also providing more detailed error explanations.
All our experiments were conducted using only quantized open-weight models and two GPUs with 48GB of VRAM, ensuring that the results can be reproduced in many AI research labs. The source code and data are available on Github. [^1] https://github.com/ivankartac/OpeNLGauge
2 Related Work
Metric No Ref. Aspects Open Err-Span
G-Eval
Prometheus
Auto-J
InstructScore
TIGERScore
Themis
[…]
Table 1: Comparison of properties in popular LLM-based metrics for NLG: reference-free (No Ref.), customizable aspects (Aspects), evaluation and data generation exclusively using open-weight LLMs (Open), and producing precise explanations with error-span annotation (Err-Span). The compared metrics are: G-Eval Liu et al. (2023), Prometheus (Kim et al., 2024b), Auto-J (Li et al., 2023a), InstructScore (Xu et al., 2023b), TIGERScore (Jiang et al., 2023), Themis (Hu et al., 2024b) and OpeNLGauge (this work).
Context: [SUMMARY OF THE STORY] Question: What is Dr. Heywood Floyd’s mission on the Clavius Base? Answer: According to the summary provided, Dr. Heywood Floyd’s mission on the Clavius Base is to investigate a recently found artifact buried four million years ago. Specifically, his mission is to ride in a Moonbus to the artifact, which is a monolith identical to the one encountered by the man-apes in the summary. Evaluation Aspect: Conciseness – Extent to which the answer is concise and to the point Explanation: This introductory phrase is an error affecting conciseness as it adds unnecessary words to the answer. The response could directly state Floyd’s mission without referencing the source (summary), given the context of the question. Severity: 2 Explanation: This phrase is considered an error affecting conciseness because it provides additional, unnecessary details about the mission. The initial sentence already clearly states Floyd’s mission, making the subsequent elaboration redundant and wordy. Severity: 3 Overall Score: Fair (3/5) Explanation of the Score: While the generated answer generally addresses the question, its conciseness is compromised by the inclusion of redundant information and an unnecessary introductory phrase. Removing these elements would enhance the answer’s directness and efficiency, potentially elevating the score to ‘Good’.
Figure 2: Example error span annotation provided by OpeNLGauge for the narrative question answering task. The answer to the question, grounded in the story summary, is evaluated for conciseness.
Although NLG has traditionally been evaluated using simple word-overlap-based metrics such as BLEU, these are known to have low correlations with human judgments Novikova et al. (2017a); Reiter (2018). This improved somewhat with the use of trained models for metrics in the past few years Yuan et al. (2021); Zhong et al. (2022); Mehri and Eskenazi (2020b), but the correlations remained moderate. Recently, numerous studies explored the application of LLMs in NLG evaluation. A prominent line of research focuses on leveraging proprietary LLMs such as GPT-4 OpenAI et al. (2024), with direct prompting for an overall score Fu et al. (2024); Kocmi and Federmann (2023b) or even annotating error spans with categories, which has been explored in machine translation Kocmi and Federmann (2023a); Fernandes et al. (2023); Lu et al. (2023). However, using proprietary models is costly and comes with a reproducibility penalty, as some LLM versions become unavailable or are modified in a non-transparent way (Chen et al., 2024). Another concern is data leakage, which affects results, but cannot be controlled in the case of proprietary models Balloccu et al. (2024).
Many recent LLM-based metrics, such as Themis (Hu et al., 2024b), TIGERScore (Jiang et al., 2023), InstructScore Xu et al. (2023b) or Prometheus Kim et al. (2024b), are built on open LLMs, but in fact they all rely on proprietary LLMs such as GPT-4 OpenAI et al. (2024) to generate, filter or annotate training data. Therefore, some of these metrics can be viewed as sophisticated knowledge distillation methods from proprietary to open-weight LLMs. This retains the reproducibility disadvantage, as reconstructing the metric from scratch or adapting it to a new task requires access to closed-source LLMs.
Although several metrics based on open-weight LLMs provide some level of interpretability or explainability, this is often limited to a short, free text review of the output text Hu et al. (2024b); Kim et al. (2024b). However, fine-grained error-span annotation offers several advantages over singular scores and comments. First, it can be easily processed automatically, allowing it to be used in post-processing steps or to provide feedback to the model or training algorithms. Second, it offers greater precision and clarity, as errors are directly associated with a particular part of the output, making it easier to identify and correct issues. Finally, it is more human-like, resembling the output of human annotation schemes such as MQM (Freitag et al., 2021) or ESA (Kocmi et al., 2024). This alignment with human evaluation makes it easier to compare metric-based assessments with human annotations and even allows the metric to serve as an initial annotator, helping to accelerate the human annotation process. While some metrics do provide annotation on the error level, these tend to use closed LLMs and their scope is generally limited to single NLG tasks, such as the machine translation approaches mentioned above or Kasner and Dusek (2024)’s work on data-to-text generation.
3 Problem Statement
We formulate the problem of evaluating the output of NLG system while providing error-based explanations as follows: Given an input $x$, an output $y$, and an evaluation aspect $a$, the task is to return a tuple $\langle z,\{e_{1},...,e_{n}\}\rangle$, where $z$ is a numeric score assigned to $y$, and $\{e_{1},...,e_{n}\}$ represents a set of error annotations. Each error annotation $e_{i}=\langle s_{i},t_{i},l_{i}\rangle$ includes a span of text $s_{i}\in y$ corresponding to the problematic segment, a textual explanation $t_{i}$, and a severity level $l_{i}$. Although the term error is used throughout this work, it should be understood in a broader sense as any issue in the text related to the evaluated aspect $a$. Examples of outputs provided by OpeNLGauge are presented in Figure 2 and in Appendix H.
4 Open LLM Ensemble as Evaluator
To achieve a high-quality evaluation of NLG output, we propose a two-stage ensemble of open-weight LLMs. The ensemble consists of $n$ annotator models, which perform independent analyses of the provided NLG output, and a supervisor model, which is responsible for merging their results and filtering inaccuracies. The overview of the approach is presented in Figure 1.
Although our approach requires multiple LLMs, using the ensemble with a handful of LLMs ($n=5$ in our experiments) is still less computationally demanding than running multiple generations to obtain statistical estimates, as required by some metrics, e.g. G-Eval ($n=20$, Liu et al. (2023)). Furthermore, we only use quantized LLMs to limit the computational requirements.
Annotator models
The annotator models are open-weight LLMs that are prompted to identify error spans in the text and to provide detailed explanations as well as severity levels for each error. This approach not only facilitates the interpretability of the evaluation process, but also acts as a chain-of-thought mechanism Wei et al. (2022), helping the model to ground its decisions in a structured reasoning path.
The full annotator model prompt is provided in Appendix E. It contains the description of the evaluated task (e.g., data-to-text), the definition of the evaluated aspect, and a template for the model’s response. Since LLMs are known to confuse different evaluation aspects (Hu et al., 2024a), the prompt also contains several rules that instruct the model to remain focused on the specific evaluation aspect, not to make additional assumptions, and to justify any score lower than the maximum by at least one identified error. Inspired by Liu et al. (2023), we also include detailed steps for error identification.
Furthermore, we provide a description of the overall scoring scale, including an explanation of the lowest and highest scores. The scale is presented as categorical (Unacceptable, Poor, Fair, Good, Excellent), based on the intuition that adjectival categorical scales may be easier for language models to interpret than numerical scales. We use a numerical severity scale (1-5) for scoring individual errors in order to avoid confusion with the overall scoring scale. [^2] A categorical scale was also initially used for error severity. However, we found that LLMs sometimes confused the error severity scale with the overall score scale.
Finally, the prompt contains the input that was used to generate the evaluated output. Some tasks may involve multiple inputs; for instance, evaluating knowledge-grounded question answering requires both question and context knowledge. In such cases, the context is also presented to the model under separate headers. Although structured formats like JSON might make parsing of the output easier, prompting LLMs to reason within strict structured outputs has been shown to impair their performance (Tam et al., 2024; Beurer-Kellner et al., 2024). Therefore, the models are instructed to produce textual outputs.
Supervisor model
The final score of the NLG output is computed by OpeNLGaugeens as a simple average of the scores provided by the annotator models. However, to meet the requirement of explainable output, the error analyses of multiple annotators need to be unified. This is the task of the supervised LMM model.
This open-weight LLM is instructed to: (1) merge errors detected by multiple models, (2) unify their output format, and (3) clean up the annotations. Specifically, the model is instructed to merge all error annotations that refer to the same issue at approximately the same location in the text, while maintaining annotation granularity. It is also prompted to fix potential deviations from the expected output format and to remove any markdown formatting added by the annotators. Finally, the supervisor should clean up outputs of annotator models that occasionally include two non-adjacent spans under a single annotation, or insert comments about the span directly into the location line. To simplify the cleaning process, error analyses produced by outlier annotator models [^3] The annotation is considered to be an outlier if the score provided by the annotator differs from the mean score by at least two standard deviations, and this difference is at least 1. are filtered out from the input to the supervisor model. The full supervisor model prompt is provided in Appendix E.
5 Training a Distilled Model
5.1 Synthetic data generation
To distill knowledge from the proposed model, we collected the outputs of over a hundred NLG systems and applied OpeNLGaugeens to produce synthetic evaluation data. We include NLG outputs produced on 15 datasets covering five task categories and almost 40 aspects. Table 2 shows the basic statistics of the constructed dataset.
The NLG outputs included cover the following tasks: data-to-text, summarization, question answering, dialogue response generation and story generation. These were chosen to represent a diverse range of tasks, each with unique objectives, input-output relationships and evaluation aspects. Data-to-text represents tasks where input is transformed from one representation to another. Summarization compresses information from a longer text into a shorter version. Question answering extracts relevant information from a context in response to a specific query. Finally, dialogue response generation and story generation create outputs with novel content. Note that the underlying datasets are used only as input for NLG systems and do not include human evaluations of any kind.
For each task, we select a set of relevant evaluation aspects with their definitions. We consider two aspects distinct if they are associated with different tasks, even if they share the same name. For instance, coherence in dialogue response generation refers to the coherence of the response with the dialogue history, while in summarization, it refers to the internal coherence of a summary itself. We list aspects and datasets for each task in Appendix A.
To obtain outputs with varying quality and diverse types of errors, we sample outputs from a variety of systems, ranging from rule-based approaches to the latest state-of-the-art LLMs. For older systems, we use pre-generated outputs from existing datasets, while outputs from more recent systems including LLMs are newly generated. An overview of the evaluated systems is shown in Appendix B.
To limit computational requirements for dataset generation, we decided to apply sampling that ensures data diversity and broad coverage of different NLG systems and aspects while significantly reducing the size of the dataset. For each input, we randomly sample the outputs of $N$ systems, and then $M$ aspects are sampled for each input-output pair. This results in $N\,\times\,M$ (input, output, aspect) triples for each input, which are then annotated by OpeNLGaugeens to obtain synthetic annotations. For most tasks, we set $N=4$ and $M$ = 3. This sampling strategy aims for a balanced distribution of inputs, NLG system outputs and evaluation aspects. Additionally, it ensures exposure to different outputs for the same input, i.e., it should not prime models to evaluate based solely on patterns in the inputs. Finally, by presenting multiple aspects for the same input-output pair, models are encouraged to learn differences in output quality across various evaluation aspects.
Task Src. Sys. Asp. Examples
Summarization 5 39 6 12,070
Data-to-text 4 29 7 07,894
Dialogue 3 36 9 10,074
Story Generation 1 9 6 03,200
Question Answering 2 15 11 04,849
Table 2: Dataset statistics: number of source datasets (Src.), systems (Sys.), evaluation aspects (Asp.) and training examples for each task.
5.2 Fine-tuning procedure
We use the dataset described in Section 5.1 for supervised fine-tuning of a specialized LLM evaluator, with an instruction-tuned version of Llama 3.1 8B as the backbone. To avoid training the LLM to predict floating-point scores, we convert them to integers in the range 0-100 and then bin them to the nearest multiple of five. This extends the output space from five to twenty values (instead of 100), which is a trade-off between greater granularity in predictions and manageable task complexity.
As the model is expected to learn the task from training data, we used a special prompt template for fine-tuning. The prompt contains only a brief description of the task, the definition of the evaluated aspect, as well as the input and output to be evaluated. The prompt is provided in the Appendix E.
Metric QAGS-CNN/DM QAGS-XSUM Average
$r$ $\rho$ $\tau$ $r$ $\rho$ $\tau$ $r$ $\rho$ $\tau$
ROUGE-1 0.338 0.318 0.248 -0.008 -0.049 -0.040 0.165 0.134 0.104
ROUGE-2 0.459 0.418 0.333 0.097 0.083 0.068 0.278 0.250 0.200
ROUGE-L 0.357 0.324 0.254 0.024 -0.011 -0.009 0.190 0.156 0.122
BERTScore 0.576 0.505 0.399 0.024 0.008 0.006 0.300 0.256 0.202
MoverScore 0.414 0.347 0.271 0.054 0.044 0.036 0.234 0.195 0.153
FactCC 0.416 0.484 0.376 0.297 0.259 0.212 0.356 0.371 0.294
QAGS 0.545 - - 0.175 - - 0.375 - -
BARTScore 0.732 0.680 0.555 0.175 0.171 0.139 0.454 0.425 0.347
UniEval 0.682 0.662 0.532 0.461 0.488 0.399 0.572 0.575 0.466
G-Eval (GPT-3.5) 0.477 0.516 0.410 0.211 0.406 0.343 0.344 0.461 0.377
G-Eval (GPT-4) 0.631 0.685 0.591 0.558 0.537 0.472 0.595 0.611 0.532
GPT-3.5 0.454 0.514 0.417 0.279 0.348 0.295 0.366 0.431 0.356
GPT-4 0.735 0.746 0.626 0.541 0.528 0.439 0.638 0.637 0.532
Llama 3.1 8B 0.275 0.242 0.219 0.218 0.230 0.218 0.247 0.236 0.219
Auto-J 0.291 0.238 0.214 0.225 0.214 0.203 0.258 0.226 0.209
TIGERScore 0.574 0.562 0.479 0.424 0.445 0.412 0.499 0.504 0.446
InstructScore 0.287 0.278 0.233 -0.096 -0.134 -0.119 0.095 0.072 0.057
Themis 0.747 0.761 0.680 0.599 0.607 0.546 0.673 0.684 0.613
OpeNLGaugeens 0.738 0.753 0.627 0.630 0.624 0.531 0.684 0.689 0.579
- Command R+ 104B 0.676 0.675 0.617 0.540 0.541 0.515 0.608 0.608 0.566
- Gemma 2 27B 0.579 0.646 0.579 0.592 0.614 0.563 0.585 0.630 0.571
- Llama 3.1 Nemotron 70B 0.705 0.733 0.650 0.587 0.586 0.540 0.646 0.659 0.595
- Mistral Large 2 123B 0.658 0.704 0.635 0.577 0.570 0.541 0.617 0.637 0.588
- Qwen 2.5 70B 0.678 0.720 0.635 0.568 0.569 0.526 0.623 0.644 0.581
OpeNLGaugeft 0.668 0.695 0.584 0.607 0.607 0.524 0.638 0.651 0.554
Table 3: Segment-level Pearson ($r$), Spearman ($\rho$) and Kendall ($\tau$) correlations of different metrics for factual consistency on QAGS. The best correlations are highlighted in bold, the second best are underlined.
6 Experimental Setup
6.1 Models
Ensemble of LLMs
The ensemble consists of six open-weight LLMs: five annotators and one supervisor. Each selected LLM is distributed under a license that permits at least non-commercial use and allows the model’s outputs to be used as training data. At the time of performing the experiments, these models ranked among the top-performing open-weight LLMs on the Chatbot Arena Leaderboard [^4] https://lmarena.ai/?leaderboard (Chiang et al., 2024).
The annotator models include the following: Llama 3.1 Nemotron 70B (Wang et al., 2024), Qwen 2.5 70B (Yang et al., 2024), Gemma 2 27B (Gemma Team et al., 2024), Command R+ 104B (Cohere For AI, 2024), and Mistral Large 2 123B [^5] https://mistral.ai/news/mistral-large-2407/. We apply Llama 3.3 70B (Dubey et al., 2024) as the supervisor model. To address computational constraints, we use quantized versions available through the Ollama platform. [^6] https://ollama.com/ For synthetic data generation, we set the temperature to zero to obtain consistent results. For details, see Appendix C.
Fine-tuning parameters
We use LoRA with rank 16 and alpha 32 to fine-tune Llama 3.1 8B. The model is trained for one epoch with a learning rate of 2e-4, using AdamW optimizer (Loshchilov and Hutter, 2017) with a weight decay of 0.01. We apply a linear learning rate schedule with a warm-up period corresponding to the first 5% of training steps. The batch size is set to 16. This setup enables resource-efficient training that requires only around 20GB of VRAM, completing one training epoch in just six hours on a single A40 GPU.
6.2 Evaluation Datasets
We used seven popular meta-evaluation datasets to assess the correlation of our approach with human judgements (see their detailed description in Appendix D). The datasets cover the following tasks: summarization – SummEval (Fabbri et al., 2021b), QAGS (Wang et al., 2020a); story generation – HANNA (Chhun et al., 2022); data-to-text – SFRES and SFHOT (Wen et al., 2015), text simplification – Wiki-DA (Alva-Manchego et al., 2021) and dialogue response generation – TopicalChat (Gopalakrishnan et al., 2023). Note that text simplification is a task unseen during training and TopicalChat contains an unseen aspect (groundedness). For human evaluation of error spans, we used the RotoWire dataset (Thomson and Reiter, 2020) containing annotations from a data-to-text task in the basketball domain, which was unseen during training by the evaluated metrics.
System Error span Error No span given Not an error Hallucination Total
Explanation
OpeNLGaugeens Correct 218 9 3 0 230
Partially correct 41 0 0 1 42
Incomplete 20 0 1 0 21
Vague 4 2 1 0 7
Incorrect 20 3 32 1 56
Not an error 0 0 10 0 10
Total 303 14 47 2 366
Span OK (%) 83
Exp. correct (%) 63
OpeNLGaugeft Correct 108 3 4 4 119
Partially correct 60 1 2 0 63
Incomplete 9 0 0 0 9
Vague 3 9 0 0 12
Incorrect 41 7 42 6 96
Not an error 0 1 8 0 9
Total 221 21 56 10 308
Span OK (%) 72
Exp. correct (%) 39
TigerScore Correct 8 17 0 0 25
Partially correct 5 2 0 0 7
Incomplete 2 2 0 0 4
Vague 0 3 0 0 3
Incorrect 5 16 2 0 23
Not an error 1 2 0 0 3
Total 21 42 2 0 65
Span OK (%) 32
Exp. correct (%) 38
Table 4: Results of human evaluation of error-span explanations. Each table shows absolute occurrence counts of different error span and explanation validity (see Section 7.2), with overall proportions of correct annotation given below.
6.3 Baselines
We compare our methods to a variety of commonly used evaluation metrics, including traditional metrics like ROUGE (Lin, 2004b), BLEU (Papineni et al., 2002b) and METEOR (Agarwal and Lavie, 2008), distance-based metrics like MoverScore (Zhao et al., 2019), BERTScore (Zhang et al., 2019), and trained metrics like BARTScore (Yuan et al., 2021) and UniEval (Zhong et al., 2022). We also include LLM-based metrics like GPTScore (Fu et al., 2024), G-Eval (Liu et al., 2023), Prometheus (Kim et al., 2024b), Auto-J (Li et al., 2023a), InstructScore (Xu et al., 2023b), TIGERScore (Jiang et al., 2023) and Themis (Hu et al., 2024b). Additionally, we report the performance of LLMs directly prompted to evaluate the outputs: GPT-3.5 and GPT-4 (see Chiang and Lee (2023b)). To measure the improvement of OpeNLGaugeft relative to the base model, we also include Llama 3.1 8B as an additional baseline. The specific metrics reported differ depending on the evaluated dataset and NLG task.
Additionally, our comparisons include some task-specific and aspect-specific metrics: QAGS (Wang et al., 2020a) and FactCC (Kryscinski et al., 2020) for evaluating factual consistency in summarization; SARI (Xu et al., 2016) and LENS (Maddela et al., 2022) for text simplification, and USR (Mehri and Eskenazi, 2020b) for dialogue response generation tasks. The latter metric has several variants; we use the variant with the best Pearson correlation for each aspect in our comparison.
7 Results
7.1 Score Correlation with Humans
Figure 3: Results of ablation experiments on QAGS and TopicalChat for OpeNLGaugeens. The plotted values represent differences in Spearman’s $\rho$ correlations with human scores between the ensemble with the original prompt and the corresponding ablation. For TopicalChat, Coh. = coherence, Eng. = engagingness, Gro. = groundedness, Nat. = naturalness, Avg. = average for all aspects.
The results for factual consistency on QAGS are presented in Table 3 and the results for other datasets are available in Appendix F. On QAGS, OpeNLGaugeens achieves the highest average performance on both Pearson and Spearman correlations, while on Kendall it is outperformed by Themis. The distilled version of our metric was consistently the third best measure. Notably, it outperformed the proprietary GPT-4 on this task.
On the other summarization benchmark, SummEval, the best performing metric was Themis, followed by OpeNLGaugeens. Comparing our approach to another method that provides a similar level of explainability (error span annotations), TigerScore, we observe 15 p.p. improvements on the Pearson correlation averaged over all aspects. The smaller OpeNLGaugeft outperformed all other fine-tuned LLM-based metrics except Themis, and also outperformed GPT-3.5 and the version of G-Eval that used it by a large margin.
On data-to-text problems, OpeNLGaugeens excelled in evaluating naturalness of generated texts, while OpeNLGaugeft showed as the strongest evaluator of informativeness on the SFRES dataset. Interestingly, on data-to-text problems, the distilled version of our model achieved slightly better results on average than our ensemble. A similar situation is observed on the story generation dataset, where the distilled version obtained better average scores on Spearman and Kendal correlations, and closely followed OpeNLGaugeens on Pearson.
Generalization to unseen aspects
On TopicalChat, the proprietary GPT-4 emerged as the strongest evaluator, with our OpeNLGaugeens performing between that model and GPT-3.5. OpeNLGaugeft achieved almost 25 p.p. improvement in both Pearson and Spearman correlations compared to its Llama 3.1 backbone. Most importantly, a large improvement is also observed on the aspect not seen during training (groudedness), where the model outperformed all fine-tuned LLM-based metrics except Themis on Pearson correlation.
Generalization to unseen tasks
For text simplification, our ensemble shows superior performance across the board, outperforming LENS, a strong baseline metric specialized for this task. OpeNLGaugeft, although not originally trained on the task, outperforms all baseline metrics on average correlation and achieves a particularly high score on meaning preservation.
7.2 Human Evaluation of Error Spans
We performed a small in-house human evaluation study to compare the quality of explanations obtained by OpeNLGauge and TigerScore, another LLM-based metric that also provides error-span annotations (see Table 1). For this purpose, we used a data-to-text task in the basketball domain (RotoWire dataset (Thomson and Reiter, 2020)). Five expert annotators evaluated the output of all three systems on 40 instances (a total of 120 outputs). We asked the annotators to: (1) evaluate provided error-spans, marking them as correct error spans, spans not containing an error, hallucinated spans, and situations where no span was provided; (2) evaluate generated explanations, marking them as correct, partially correct, incomplete, vague, incorrect, or texts that do not describe an error.
The results are shown in Table 4. Both OpeNLGaugeft and OpeNLGaugeens were over two times more accurate than TigerScore at annotating error spans, while finding over ten times more correct error spans. The task of providing accurate error explanations was more difficult for all the approaches evaluated. OpeNLGaugeens achieved the highest performance and was almost twice as accurate as TigerScore and OpeNLGaugeft, which achieved similar accuracies.
7.3 Ablation Experiments
To better analyze the proposed approach, we perform ablation experiments on the selected prompt and ensemble structure.
Prompt ablations
We explore the effect of using different scales for the overall score and error severity scales: integer scales for both (Ablation 1), integer scale for overall and categorical scale for severity (Ablation 2), categorical scale for both (Ablation 3). Recall that OpeNLGauge uses a categorical scale for overall score and integer scale for severity – see Section 4.
The differences in the correlations between the full prompt and the ablations are summarized in Figure 3, for detailed results see Appendix G.1. Overall, the change of scale has little effect on the average correlation of the whole ensemble, but it has a dramatic effect on some individual LLMs and aspects. This illustrates the ability of the ensemble to compensate for weaknesses in individual models.
Finally, we examine the effect of removing the evaluation rules from the prompt (Ablation 4), which has an inconsistent effect on different models/aspects, but on average degrades the ensemble evaluation quality – up to 5.6 p.p.
Ensemble structure
We also analyze the effect of ensemble size and the influence of its particular components on the correlations with human scores by recomputing the results for all ensemble combinations. The results are presented in Appendix G.2. For Wiki-DA, performance increases with ensemble size, with the full ensemble being the best combination. For other datasets, there are a few smaller combinations that actually rank higher, but none is consistently better than the full ensemble.
8 Summary
In this work, we propose OpeNLGauge – a versatile method for evaluating NLG that uses only open-weight models and provides fine-grained explainability. The method provides a much better explanation quality than previous methods and achieves competitive correlations with human judgments.
Limitations
OpeNLGauge is a method for evaluating a variety of NLG tasks. While this paper presents the evaluation on several NLG tasks and the method achieves good performance on unseen aspects, domains and tasks, the actual performance on new NLG tasks is unknown. In particular, the metric has not been tested in a multilingual setting. Moreover, previous research has shown that some LLM-based metrics have a bias towards texts generated by LLMs Liu et al. (2023).
Acknowledgments
This work was supported by the European Research Council (Grant agreement No. 101039303, NG-NLG) and used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
Appendix A Aspects and datasets in the training data
This section provides a detailed overview of source datasets and evaluation aspects used to generate our training dataset.
A.1 Summarization
As source data for summarization, we utilize five datasets that include three distinct tasks: news article summarization, forum post summarization and dialogue summarization. Our evaluation dataset includes six commonly used aspects: four related to meaning of the evaluated text, one to its form and one to both. Some of these aspects overlap significantly in their definitions. This is intentional, as our goal is to make the evaluator model robust to small variations in aspect definitions, and enable generalization to new aspects. Table 6 provides an overview of the evaluated systems.
A.1.1 Datasets
  • CNN/DailyMail (Hermann et al., 2015) is a popular summarization dataset that consists of news articles from CNN and DailyMail, paired with corresponding bullet-point summaries. Originally developed for question answering, it was later adapted for summarization (Nallapati et al., 2016). For this dataset, we use pre-generated outputs from Stiennon et al. (2020), which include results from 11 different systems, including human references and an extractive baseline. Since our meta-evaluation datasets contain inputs from CNN/DailyMail, we ensure there is no overlap in the source texts between these datasets when sampling the inputs.
  • Newsroom (Grusky et al., 2018) is a large-scale dataset of news articles and their summaries, collected from diverse sources, domains, authors and time range. In addition to newly generated LLM outputs, we use outputs of the systems evaluated in the original paper, which include three summarization systems and two extractive baselines.
  • SAMSum dataset (Gliwa et al., 2019) addresses the dialogue summarization task, containing dialogues with short summaries created by linguists. In addition to outputs from several different LLMs, we use pre-generated outputs of six systems from (Gao and Wan, 2022).
  • TL;DR (Völske et al., 2017) is a collection of Reddit posts with user-created summaries. Unlike many popular datasets that focus on news articles, TL;DR contains informal and less structured texts spanning diverse topics. Similarly to CNN/DailyMail, we use outputs from Stiennon et al. (2020).
  • XSum (Narayan et al., 2018) is a dataset for extreme summarization task, designed for abstractive approaches. Unlike datasets such as Newsroom or CNN/DailyMail, which favor extractive summarization, XSum contains BBC articles paired with concise, single-sentence summaries. We utilize pre-generated outputs from five systems and baselines evaluated in the original paper, along with newly generated output from LLMs.
A.1.2 Aspects
  • Consistency evaluates whether the summary is factually aligned with the source text. This involves determining if the facts in the summary can be entailed by the source. Consistency is closely tied to hallucinations, which may be categorized as factual or non-factual. In our approach, all information is required to be supported by the source text, making both types of hallucinations inconsistent with the source. The definition of consistency used in our prompts is: Extent to which the facts in the summary are consistent with the source text. Factually consistent summary should not contain facts that are not supported by the source text.
  • Accuracy is largely synonymous with consistency. It evaluates whether the factual information from the source is accurately represented in the summary. Following Stiennon et al. (2020), we define accuracy as: Extent to which the factual information in the summary accurately matches the source text. An accurate summary should not contain information that is not present in the source text, should not contradict the source text, and generally should not be misleading.
  • Relevance of a summary is concerned with content selection. A relevant summary should include important points from the source text while omitting unimportant details. Compared to consistency, relevance is more subjective, as determining what information should or should not be included in a summary can sometimes be ambiguous. The definition used in our evaluation is: Extent to which the summary captures important information of the source text. A relevant summary should include all and only important information from the source text.
  • Coverage evaluates how much of the important information from the source text is covered by the summary. In this sense, it is closely related to relevance, however, it does not include non-redundancy as a criterion. We use a definition adapted from Stiennon et al. (2020): Extent to which the summary covers the important information in the source text. A summary has good coverage if it mentions the main information from the source text that is important to understand the events described in the text. A summary has poor coverage if someone reading only the summary would be missing several important pieces of information about the event in the source text.
  • Coherence refers to the structural quality of the summary and involves attributes such as cohesion, consistency and relevance (Reinhart, 1980). It is determined by both semantic and formal structure of the text. We define coherence as: Extent to which the summary is well-structured and organized, presenting information in a logical order that flows naturally from sentence to sentence. Coherent summary forms a unified body of information and makes it easy to understand the main ideas.
  • Fluency focuses on the formal quality of the text, including grammaticality and naturalness. Unlike coherence, fluency is concerned with sentence–level quality rather than the overall structure of the text. We define fluency as: Formal quality of individual sentences of the summary. A fluent sentence should be grammatical, natural and easy to understand.
Format Example Datasets
JSON {"name": "The Phoenix", "eatType": "pub", "food": "Indian", ...} E2E NLG
[…]
attribute-value (2) name[The Phoenix], eatType[restaurant], food[Indian], ... E2E NLG
RDF (1) "The_Velvet_Underground | genre | Proto-punk" ... WebNLG 2020
RDF (2) (The_Velvet_Underground, genre, Proto-punk) ... WebNLG 2020
[…]
linearized table <page_title> List of Norwegian fjords </page_title> <section_title> ... ToTTo
Table 5: Input data formats for data-to-text tasks.
A.2 Data-to-text
For data-to-text category, we collected inputs from four datasets that represent four distinct tasks: table-to-text, RDF-to-text, attribute-value list to text and logical NLG. Since it is important for an NLG evaluation method to reliably evaluate outputs with respect to structured data in different formats, we include a number of different input formats in our training data, including JSON, CSV and linearized tables with markup. Table 5 provides an overview of the input formats. The list of evaluated systems is shown in Table 7.
A.2.1 Datasets
  • E2E NLG dataset (Dušek et al., 2020; Novikova et al., 2017b) was chosen as a representative of simple data-to-text tasks. In this dataset, models are tasked with generating descriptions of restaurant venues based on attribute-value-based meaning representations (MRs). Target descriptions were crowdsourced using textual and pictorial representations of MRs as stimuli. To ensure diversity, we selected pre-generated outputs from five systems described by (Dušek et al., 2020), considering their model architectures and evaluation results across various metrics and aspects. These were extended by new outputs from several LLMs. The inputs and pre-generated outputs were sampled from the test set. Three different input formats are used in the training data (see Table 5).
  • WebNLG 2020 (Ferreira et al., 2020) is an RDF-to-text dataset designed for generating natural language text from RDF triples collected from DBPedia knowledge base (Mendes et al., 2012). Each input contains between one and seven triples, where each triple represents a binary relation in the form (subject, property, object). Similarly to E2E NLG, we selected a diverse set of output systems based on model architectures and evaluation results from the WebNLG+ 2020 Challenge, and used additional LLMs to generate new outputs.
  • ToTTo (Parikh et al., 2020) is an open-domain table-to-text generation dataset. It consists of Wikipedia tables with highlighted cells, and the task is to generate single–sentence descriptions of the data in these highlighted cells. Inputs are provided in two formats: full tables, which include indices to highlighted cells, and linearized tables, where only the highlighted data is presented in a linear order, while the structure is annotated with markup tags. As we observed that even medium-sized open-weight LLMs often struggle with hallucinations when using full tables, we restricted our evaluation to linearized tables.
  • LogicNLG (Chen et al., 2020) introduced the task of logical NLG, where models generate statements that can be logically entailed from the data in a table. This task involves various aggregations and comparisons, making it more difficult than simply transforming structured data to free-form text. Although the task explicitly requires generating five logical statements per input, we sampled between one to five statements from each generated output to increase the diversity of output lengths. Inputs were formatted as CSV with “|” as a separator. In addition to new LLM outputs, we used existing system output from Zhao et al. (2023b).
A.2.2 Aspects
  • Faithfulness measures whether all information in the generated text is supported by the data, making it equivalent to precision. Similarly to factual consistency in summarization, we consider both factual and non–factual hallucinations as errors. For our evaluations, we define faithfulness as: Extent to which the information in the text is supported by the data.
  • Correctness evaluates whether the information from the data is accurately presented in the generated text. The output is maximally correct if it does not contain any incorrect statements with respect to the input. Correctness overlaps significantly with faithfulness, but its definition varies based on the specific task. For example, we define correctness for LogicNLG as: Extent to which the statements are logically and factually correct with respect to the provided data.
  • Coverage refers to the degree to which the generated text covers the information in the data. The output has maximum coverage when all information from the data is included in the text, which makes it analogous to recall. For example, in WebNLG 2020, it evaluates whether all predicates and their arguments are mentioned, while in ToTTo, it determines whether all highlighted table cells are described. We apply coverage to evaluate all datasets except LogicNLG, where the task is to infer interesting observations, rather than fully cover the source data. We define coverage as follows, with slight variations depending on the task: Extent to which the text includes description of all information presented in the data.
  • Informativeness is closely related to coverage, as it evaluates how much of the information the generated text provides. However, it does not require complete coverage of the data, therefore it is applicable to tasks such as LogicNLG, where full coverage of a table is not necessary. The definition used in our evaluation depends on the particular dataset. For example, the definition used for LogicNLG is: Extent to which the statements provide interesting or useful information about the data.
  • Fluency refers to the formal quality of the generated text, and includes grammaticality, naturalness and readability. Some definitions also include coherence (Ferreira et al., 2020), although coherence is usually treated as a separate aspect. For our evaluation, we define fluency as: Extent to which the text is grammatical, natural and easy to understand.
  • Grammaticality focuses on the correctness of grammar and spelling in the generated text. A text is fully grammatical if it contains no grammatical or spelling errors. While grammaticality is often included as a sub-aspect of fluency, both aspects are commonly used in practice. Therefore, we include it to help the model learn differences between evaluation aspects on different levels of hierarchy. In our dataset, grammaticality is defined as: Extent to which the text is grammatical (free of grammar and spelling errors).
  • Naturalness refers either to the human-likeness of the text, or the likelihood that it was produced by a native speaker. Like grammaticality, naturalness is often treated as a component of fluency. Additionally, its evaluation often includes assessment of grammaticality, as this can often be an indicator of whether the text was produced by a native speaker. This illustrates how evaluation aspects often overlap or have hierarchical relationships. For our purposes, naturalness is defined as: Extent to which the text is likely to have been produced by a native speaker.
A.3 Dialogue Response Generation
For dialogue response generation, we source the inputs from three dialogue datasets, focusing on open-domain non-task-oriented dialogue response generation. Table 8 provides an overview of the evaluated systems.
A.3.1 Datasets
  • Wizard of Wikipedia (Dinan et al., 2018) consists of conversations grounded in one of 1365 topics and corresponding knowledge retrieved from Wikipedia. In these conversations, either participant may select the topic and initiate the discussion, although they have asymmetric roles. One participant takes on the role of the wizard, an expert with access to a topic-relevant knowledge, on which they can base their responses. The other participant acts as an apprentice, a curious learner that is eager to discuss the chosen topic. To create inputs of varying length, we randomly select a dialogue history length between two turns and the full conversation, and truncate the dialogue to this length. The last utterance is replaced by a system output, except when the reference is used as evaluated output. In addition to newly generated LLM responses, we include human responses from the original dataset in the outputs.
  • EmpatheticDialogues (Rashkin, 2018) is a dataset of dialogues grounded in emotional situations, designed to train and evaluate dialogue models on empathetic response generation. Each conversation is associated with an emotional label, where one of the participants describes a situation in which they experienced a given emotion. We sample up to five turns from each dialogue and replace the last utterance with a generated response. The emotion label is used as additional context for annotator LLMs to evaluate the appropriateness and empathy of the responses, but is excluded from the prompts used for system output generation.
  • DailyDialog (Li et al., 2017) includes conversations on various daily life topics, annotated with emotion labels and communicative intents. We include data from DailyDialog to represent diverse topics and scenarios in the training set. Alongside newly generated responses, we also collect pre-generated outputs from three sources (Gupta et al., 2019; Huang et al., 2020; Zhao et al., 2020) to represent older dialogue systems.
A.3.2 Aspects
  • Coherence in dialogue is a concept slightly different from coherence in tasks that involve generation of standalone texts, such as summarization or story generation. In dialogue, it measures how meaningful and logically consistent the response is with the preceding conversation. This includes not only alignment of the response with the last utterance, but also consistency with the dialogue participant’s earlier responses in terms of logic and style. In our evaluation, coherence is broadly defined as: Extent to which the response is a meaningful continuation of previous dialogue.
  • Relevance evaluates how closely a response aligns with the topic of conversation. In this sense, this aspect overlaps with coherence to some degree. We define relevance as: Extent to which the response is relevant and on-topic given the dialogue history.
  • Appropriateness addresses whether the response is semantically and pragmatically appropriate in the given context. Depending on the definitions, appropriateness might overlap to a large extent with coherence and relevance. For our evaluation, we define appropriateness as: Extent to which the response is semantically and pragmatically appropriate given the conversation history.
  • Empathy is evaluated specifically on responses from the EmpatheticDialogues dataset, where the goal is to determine whether the response acknowledges and reflects the emotions of the other participant. We define empathy as: Extent to which the response shows understanding of the feelings of the person talking about their experience.
  • Interestingness is concerned with the informational value of the response, specifically whether it presents stimulating ideas, facts or opinions. Although it is one of the more subjective evaluation aspects, we define it vaguely and let the annotator models determine the criteria for interestingness. Therefore, in our dataset, we define interestingness as: Extent to which the response is interesting given the dialogue history.
  • Engagingness is closely related to interestingness but is sometimes treated as a distinct aspect (e.g., Mehri and Eskenazi, 2020a; See et al., 2019). While interestingness focuses on the context itself, engagingness emphasizes maintaining the user’s attention and encouraging them to continue with the conversation. For our purposes, engagingness is defined as: Extent to which the response captures and maintains the user’s interest, encouraging further interaction. Engaging responses contain opinions, preferences, thoughts or interesting facts.
  • Fluency in dialogue response generation has a similar meaning to its use in other tasks and refers to the formal quality of the response. We define fluency as: Extent to which the response is grammatically correct, natural and fluent.
  • Understandability evaluates both the content and form of a response, focusing on its clarity and ease of comprehension. The definition we use for our evaluation is: Extent to which the response is easy to understand and comprehend given the dialogue history.
A.4 Story Generation
The NLG tasks discussed so far generally contain inputs that are relatively longer compared to the outputs. This pattern is especially common in tasks like summarization and dialogue generation, although certain data-to-text tasks also share this characteristic. To represent scenarios with short inputs and long outputs, we include story generation in the training data. Table 9 lists the evaluated systems.
A.4.1 Datasets
As the source of inputs, we use WritingPrompts (Fan et al., 2018), a story generation dataset derived from Reddit’s WritingPrompts subreddit, where users submit prompts that can inspire other users to write stories. The dataset consists of a diverse range of topics, story lengths and writing styles. We reuse existing outputs from the OpenMEVA dataset (Guan et al., 2021) and generate additional outputs by four LLMs. To increase the diversity of generated stories in terms of their length, we generate the outputs with two different prompt versions, each requiring a different length of the story. The outputs are then randomly sampled from either the shorter or the longer set. As we observed a tendency of LLMs to generate a long list of errors for longer inputs, our evaluator models are instructed to limit the number of identified errors to a maximum of eight, and to prioritize the most severe ones if necessary.
A.4.2 Aspects
While relevance and coherence are two commonly used aspects for story generation, there is no consensus on which other evaluation aspects are the most relevant. Inspired by social sciences, Chhun et al. (2022) propose four additional aspects, aimed at providing a complete and non-redundant set of criteria. Following their work, we adopt the aspects defined in the HANNA benchmark, using our own definitions for most of them:
  • Relevance measures the degree to which a story aligns with the given prompt (Chhun et al., 2022; Chiang and Lee, 2023a), title (Jhamtani and Berg-Kirkpatrick, 2020; Yao et al., 2019; Xie et al., 2023) or story beginning Wang et al. (2020b). In some cases, relevance also evaluates whether the story remains on-topic for its duration (e.g., Goldfarb-Tarrant et al., 2020). Since our inputs are prompts, we define relevance as: Extent to which the story is relevant to the writing prompt.
  • Coherence in story generation typically refers to logical consistency and narrative flow (e.g., Yao et al., 2019; Li et al., 2023b; Jhamtani and Berg-Kirkpatrick, 2020). Other works define coherence more vaguely, such as how much the story “makes sense” (Chhun et al., 2022) or how well its sentences “fit together” (Xie et al., 2023). For our purposes, coherence is defined as: Extent to which the story is logically consistent and coherent.
  • Engagement is a subjective and often vaguely defined aspect that evaluates how engaging the story is to the reader (e.g., Chhun et al., 2022; Li et al., 2023b). Due to its inherent subjectivity, we apply a simple definition and leave its interpretation to the evaluator models: Extent to which the story is engaging and interesting.
  • Empathy is related to emotional commentary and empathy, and refers to how well the story conveys character’s emotions. We define empathy as: The clarity and depth with which the character’s emotions are conveyed in the story.
  • Surprise is concerned with the story’s ending, and evaluates its unexpectedness and originality. We define surprise as: How surprising the end of the story was.
  • Complexity measures how intricate and elaborate the story is. Complexity is not necessarily an aspect of quality, but rather a feature of the text. Whether greater complexity is desired or not depends on the audience. Our definition of complexity is: How elaborate the story is.
A.5 Question Answering
The question answering subset of our dataset consists of two distinct tasks: narrative question answering and table question answering. The inputs consist of a question and structured data in which the answer should be grounded. The evaluated systems are listed in Table 10.
A.5.1 Datasets
  • NarrativeQA (Kočiský et al., 2018) consists of human-written questions and free-form answers based on stories or their summaries. The stories include books and movie scripts and are provided either as full texts or as human-written summaries. Since full stories do not fit into the context window of many LLMs, we use only summaries to generate the outputs. Along with newly generated answers, we include human reference answers from the original datasets in the evaluated outputs.
  • FeTaQA (Nan et al., 2022b) is a question answering dataset based on Wikipedia tables that requires models to aggregate and reason about the entities in the table and their relations. This places the task at the intersection of data-to-text and question answering task categories. In addition to new LLM outputs, we use pre-generated outputs from Zhao et al. (2023b).
A.5.2 Aspects
  • Correctness evaluates if the answer to a question is correct with respect to the input. Since our models are instructed to assess the quality on an ordinal scale, we evaluate a degree of correctness - the answer should receive the maximum score if it is fully correct, while lower scores should reflect the number and severity of correctness issues. We define correctness as: Extent to which the answer to the question is correct with respect to the input.
  • Informativeness addresses whether all information required by the question is provided in the answer. We define informativeness as: Extent to which the answer provides all information that the question asked for.
  • Completeness evaluates comprehensiveness of the answer and the degree to which all aspects of the question are covered. The meaning is slightly different from informativeness, which is concerned with the information that the question explicitly asks for. In our dataset, completeness is defined as: Extent to which the answer is comprehensive and ensures all question aspects are addressed.
  • Conciseness measures the degree to which an answer is focused and directly answers the question without unnecessary details and elaboration. Although the goal might often be to generate both complete and concise answers, these two aspects might correlate negatively. Conciseness is defined as: Extent to which the answer is concise and to the point.
  • Relevance is concerned with specificity of the answer and measures the degree to which the answer addresses the particular question asked. Although it is related to conciseness, relevance is not that much concerned with the amount of detail in the answer. We define relevance as: Extent to which the answer is specific and meaningful with respect to the question.
  • Factuality evaluates factual consistency of the answer with the provided context. In our dataset, this aspect is used in the narrative question answering task, and we use a similar definition as in summarization: Extent to which the answer is supported by the summary.
  • Faithfulness is used for the table question answering task and is synonymous with factuality. We define faithfulness as: Extent to which the information presented in the answer is supported by the input.
  • Fluency in question answering refers to the formal quality of the answer and is defined similarly as in the other tasks presented so far: Extent to which the response is grammatical, natural and easy to understand.
  • Naturalness is interpreted in the same way as in data-to-text and dialogue response generation tasks, and the definition we apply is: Extent to which the answer is likely to have been produced by a native speaker.
  • Grammaticality measures the grammatical quality of the answer and is defined as: Extent to which the answer is grammatical (free of grammar and spelling errors).
Appendix B Collected system outputs
Tables 610 list the evaluated systems for each task category in our dataset.
Table 6: Overview of the evaluated systems in the dataset for the summarization task. Sources: GW = Gao and Wan (2022), GR = Grusky et al. (2018), NA = Narayan et al. (2018), ST = Stiennon et al. (2020), new = newly generated.
System Type Sources
Qwen 2.5 0.5B (Yang et al., 2024) Instruction-tuned LLM new
Llama 2 7B Chat (Touvron et al., 2023) Instruction-tuned LLM new
Gemma 2 2B (Gemma Team et al., 2024) Instruction-tuned LLM new
Nous Hermes 2 Mixtral 8x7B DPO [^7] https://huggingface.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-DPO Instruction-tuned LLM new
GPT-4o [^8] https://openai.com/index/hello-gpt-4o/ Instruction-tuned LLM new
OpenAI summarization (pre-trained) pre-trained LMs ST
OpenAI summarization (supervised) LMs trained with SFT ST
OpenAI summarization (RLHF) LMs trained with SFT+PPO ST
T5 (Raffel et al., 2020) pre-trained LM ST
UniLM (Dong et al., 2019) pre-trained LM ST
CODS (Wu et al., 2021) BART-based hybrid model GW
ConvoSumm (Fabbri et al., 2021a) BART-based model GW
Ctrl-DiaSumm (Liu and Chen, 2021) BART-based model GW
ConvS2S (Gehring et al., 2017) Convolutional seq-to-seq NA
Topic-ConvS2S (Narayan et al., 2018) Topic-conditioned ConvS2S NA
S2S (Cho et al., 2014; Sutskever et al., 2014) RNN-based seq-to-seq with attention GR
PNG (Vinyals et al., 2015; Gulçehre et al., 2016) Pointer-generator network GR, GW, NA
TextRank (Barrios et al., 2016) Ranking-based extractive summarization GR
Reference Human-written reference ST
Title Extractive baseline ST
LEAD Extractive baseline NA
LEAD-2 Extractive baseline ST
LEAD-3 (See et al., 2017) Extractive baseline GR, GW, ST
Ext-Oracle Extractive oracle NA
Fragments Extractive oracle NA
Table 7: Overview of the evaluated systems in the dataset for the data-to-text task. Sources: DU = Dušek et al. (2020), FE = Ferreira et al. (2020), ZH = Zhao et al. (2023b), new = newly generated.
System Type Sources
Qwen 2.5 Coder 1.5B (Yang et al., 2024) Instruction-tuned LLM new
Gemma 2 2B (Gemma Team et al., 2024) Instruction-tuned LLM new
Llama 2 7B Chat (Touvron et al., 2023) Instruction-tuned LLM new
Solar 10.7B (Kim et al., 2024a) Instruction-tuned LLM new
DeepSeek Coder v2 16B (Zhu et al., 2024) Instruction-tuned LLM new
Nous Hermes 2 Mixtral 8x7B DPO [^9] https://huggingface.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-DPO Instruction-tuned LLM new
Claude 3.5 Sonnet [^10] https://www.anthropic.com/news/claude-3-5-sonnet Instruction-tuned LLM new
GPT-4o [^11] https://openai.com/index/hello-gpt-4o/ Instruction-tuned LLM new
NILC (Sobrevilla Cabezudo and Pardo, 2020) Fine-tuned BART FE
Orange-NLG (Montella et al., 2020) Fine-tuned BART FE
Amazaon AI (Shanghai) (Guo et al., 2020) Graph CNN + T5 FE
GPT2-C2F (Chen et al., 2020) Fine-tuned GPT-2 ZH
LoFT (Zhao et al., 2023a) Fine-tuned BART ZH
PLOG (Liu et al., 2022a) Fine-tuned T5 ZH
R2D2 (Nan et al., 2022a) Fine-tuned T5 ZH
Flan-T5 (Chung et al., 2024) Instruction-tuned LM ZH
Adapt (Elder et al., 2018) RNN seq-to-seq DU
Sheff2 (Chen et al., 2018) RNN seq-to-seq DU
Slug (Juraska et al., 2018) RNN + convolutional seq-to-seq DU
Forge1 (Mille and Dasiopoulou, 2018) Rule-based DU
TR2 (Smiley et al., 2018) Template-based DU
DANGNT-SGU (Tran and Nguyen, 2020) Template-based FE
RALI-Université de Montréal (Lapalme, 2020) Template-based FE
Table 8: Overview of the evaluated systems in the dataset for the dialogue response generation task. Sources: GU = Gupta et al. (2019), HU = Huang et al. (2020), ZH = Zhao et al. (2020), new = newly generated.
System Type Sources
Claude 3.5 Sonnet [^12] https://www.anthropic.com/news/claude-3-5-sonnet Instruction-tuned LLM new
GPT-4o [^13] https://openai.com/index/hello-gpt-4o/ Instruction-tuned LLM new
Tülu 3 (Lambert et al., 2024) Instruction-tuned LLM new
Dolphin 2.9 Llama 3 8B [^14] https://huggingface.co/cognitivecomputations/dolphin-2.9-llama3-8b Instruction-tuned LLM new
Vicuna 7B (Zheng et al., 2023) Instruction-tuned LLM new
BlenderBot-small (Roller et al., 2021) Dialogue LM new
DialoGPT-small (Zhang et al., 2020) Dialogue LM new
GPT-Neo 125M (Gao et al., 2020) Pre-trained LM new
GPT-2 (Wolf et al., 2019) Pre-trained LM ZH
CVAE (Zhao et al., 2017) Conditional variational autoencoder GU
HRED (Serban et al., 2016) Hierarchical recurrent encoder-decoder GU, ZH
Transformer-generator (Dinan et al., 2018) Transformer-based generative model HU
Transformer-ranker (Urbanek et al., 2019) Transformer-based ranking model HU
DualEncoder (Lowe et al., 2015) LSTM dual encoder GU
VHRED (Serban et al., 2017) Latent variable HRED ZH
S2S (Cho et al., 2014; Sutskever et al., 2014) RNN-based seq-to-seq with attention GU, ZH
Table 9: Overview of the evaluated systems in the dataset for the story generation task. Sources: GU = Guan et al. (2021), new = newly generated.
System Type Sources
Gemma 2 2B (Gemma Team et al., 2024) Instruction-tuned LLM new
Dolphin 2.9 Llama 3 8B [^15] https://huggingface.co/cognitivecomputations/dolphin-2.9-llama3-8b Instruction-tuned LLM new
Nous Hermes 2 Mixtral 8x7B DPO [^16] https://huggingface.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-DPO Instruction-tuned LLM new
GPT-4o [^17] https://openai.com/index/hello-gpt-4o/ Instruction-tuned LLM new
GPT-2 (Radford et al., 2019) Pre-trained LM GU
GPT-KG (Guan et al., 2020) Knowledge-enhanced GPT-2 GU
Fusion (Fan et al., 2018) Convolutional seq-to-seq with attention GU
Plan&Write (Yao et al., 2019) Hierarchical RNN-based model GU
S2S (Cho et al., 2014; Sutskever et al., 2014) RNN-based seq-to-seq GU
Table 10: Overview of the evaluated systems in the dataset for the question answering task. Sources: KO = Kočiský et al. (2018), ZH = Zhao et al. (2023b), new = newly generated.
System Type Sources
Claude 3.5 Sonnet [^18] https://www.anthropic.com/news/claude-3-5-sonnet Instruction-tuned LLM new
GPT-4o [^19] https://openai.com/index/hello-gpt-4o/ Instruction-tuned LLM new
Nous Hermes 2 Mixtral 8x7B DPO [^20] https://huggingface.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-DPO Instruction-tuned LLM new
DeepSeek Coder v2 16B (Zhu et al., 2024) Instruction-tuned LLM new
Llama 2 7B Chat (Touvron et al., 2023) Instruction-tuned LLM new
Qwen 2.5 Coder 1.5B (Yang et al., 2024) Instruction-tuned LLM new
Qwen 2.5 0.5B (Yang et al., 2024) Instruction-tuned LLM new
GPT-Neo 125M (Gao et al., 2020) Pre-trained LM new
BART (Lewis, 2019) Pre-trained LM ZH
Flan-T5 (Chung et al., 2024) Instruction-tuned LM ZH
OmniTab (Jiang et al., 2022) Table pre-trained LM ZH
ReasTAP (Zhao et al., 2022) Table pre-trained LM ZH
TAPEX (Liu et al., 2022b) Table pre-trained LM ZH
Reference Human-written reference KO
Appendix C LLMs used in the experiments
The specific versions of LLMs used in the experiments are presented in Table 11.
Model Quantization Tag
Command R+ 104B 5-bit command-r-plus:104b-08-2024-q5_K_M
Gemma 2 27B 8-bit gemma2:27b-instruct-q8_0
Llama 3.1 Nemotron 70B 8-bit nemotron:70b-instruct-q8_0
Mistral Large 2 123B 4-bit mistral-large:123b-instruct-2407-q4_K_M
Qwen 2.5 70B 8-bit qwen2.5:72b-instruct-q8_0
Table 11: Quantization levels and Ollama tags used for the models.
Appendix D Meta-evaluation datasets
We used the following datasets for meta-evaluation:
  • SummEval (Fabbri et al., 2021b) is a standard meta-evaluation dataset for summarization. It consists of summaries generated from CNN/DailyMail articles, with 100 input articles and 16 different system outputs for each article. Human evaluations address four aspects: consistency, relevance, coherence, and fluency. Each output is scored by three expert annotators and five crowdworkers. We use only averaged expert annotations.
  • QAGS (Wang et al., 2020a) contains annotations of consistency for summaries from CNN/DailyMail and XSum datasets. It includes 235 CNN/DailyMail summaries and 239 XSum summaries, each annotated by three evaluators. Annotators assign a binary factual consistency score (yes/no) for each sentence of the summary. We follow (Wang et al., 2020a) and apply majority voting for each sentence annotation, followed by averaging sentence-level scores to obtain the overall score for the summary.
  • HANNA (Chhun et al., 2022) is a benchmark for story generation based on the WritingPrompts dataset. It contains human evaluations for 1056 stories across six aspects: relevance, coherence, engagement, empathy, surprise and complexity (see Appendix A.4 for details).
  • SFRES and SFHOT (Wen et al., 2015) are used to perform a meta-evaluation for data-to-text. These datasets consist of dialogue acts (DAs) in structured format with generated responses providing information about restaurants and hotels in San Francisco. Human judgments are provided for informativeness and naturalness.
  • TopicalChat (Gopalakrishnan et al., 2023) annotations from the USR dataset (Mehri and Eskenazi, 2020b) are used for meta-evaluation on dialogue tasks. The dataset includes human evaluations for five aspects: groundedness, coherence, interestingness, naturalness and understandability. As the source data differ from our training data, TopicalChat serves as out-of-domain evaluation dataset. Additionally, since groundedness is not present in our training data, we evaluate it as an unseen aspect.
  • Wiki-DA (Alva-Manchego et al., 2021) is a dataset of DA human ratings for text simplification, an unseen task for OpeNLGaugeft during training. Along with scores for fluency, this dataset also includes two unseen aspects: meaning preservation and simplicity.
Appendix E Prompt Templates
Prompt templates for annotator and supervisor LLMs are presented in Listings 12. Prompt template used for fine-tuning and inference of the distilled model is shown in Listing 3.
Appendix F Full Results
Tables 1218 contain the meta-evaluation results for additional datasets described in Appendix D.
Appendix G Ablation experiments
G.1 Prompt ablations
The detailed results of prompt ablation experiments are presented in Figure 1013.
G.2 Ensemble size ablations
The effect of the ensemble size on Spearman’s correlations are presented in Figures 49.
Appendix H Output Examples
Figures 1416 show additional output examples for RDF-to-text, dialogue response generation, and summarization tasks.
### Instructions
Your task is to evaluate an output of data-to-text task, where the model was instructed to write a single-paragraph description of a venue based on the given data. The data consist of a set of attribute-value pairs in the form attribute: value’.
Based on the given data and the generated text, identify errors in the text with respect to {{ aspect_name }} (described below).
For each error, determine its severity on a scale from 1 to 5, where 1 is the least severe and 5 is the most severe.
Definition of {{ aspect_name }}:
{{ aspect_definition }}
Rules:
Do not make assumptions and do not bring in external knowledge not present in the provided context.
Identify only the errors related to the {{ aspect_name }} of the text. Do not consider other aspects like {{ negative_aspects }}!
If there are no errors related to {{ aspect_name }} in the text, you should output No Error and provide Excellent score.
Steps:
1. Carefully read the data and identify the main attributes and their values.
2. Read the generated text and compare it with the source data with respect to {{ aspect_name }}.
3. If the text contains any error that negatively affects its {{ aspect_name }}, identify its exact location (specific word or phrase), explain why it is considered an error, and determine the severity of the error.
4. Finally, provide an overall score for the {{ aspect_name }} of the text. The score should be a label on the following scale (lowest to highest): Unacceptable’, Poor’, Fair’, Good’, Excellent’. The score Unacceptable indicates that the text is {{ min_score_desc }}, while Excellent indicates that the text is {{ max_score_desc }}.
### Data
{{ input }}
### Generated Text
{{ output }}
### Output format:
Generate your output exactly in this format:
Error 1:
Location: <location of the error - the exact word or phrase in the response>
Explanation: <explanation for the error, including the reason why it is considered {{ aspect_name }} issue>
Severity: <integer from 1 to 5>
Error 2:
Overall score: <one of: Unacceptable, Poor, Fair, Good, Excellent>
Explanation of the score: <explanation of the score>
Listing 1: Annotator prompt template for data-to-text task
### Instructions
You are given multiple error annotation sets for an AI model output. Your task is to merge the annotation sets to a single final annotation set.
The result shouldnt contain any duplicates. If there are multiple error annotations for approximately the same location that describe the same issue, you should merge them into single location. Otherwise, the error annotations should be as granular as possible. If there are multiple different locations with the same issue, each should have its own error annotation. Likewise, if there are multiple issues with respect to the same location, each should have its own error annotation. Use the following guidelines:
* Each error annotation should describe a single issue.
* Merge only annotations where the locations have significant overlap.
* When merging multiple locations, choose a single span from the output text that covers the locations from merged annotations.
* Never include multiple spans from the annotations under the same "Location" line.
* Do not include any other text in "Location" line than the text that is actually in the output, except for annotations that mention omissions or similar issues.
* When merging explanations, combine the most relevant information from the merged annotations.
* Severity levels range from 1 to 5 from least severe to most severe. Use the most severe level from the merged annotations.
* Final annotation set should not include more than 8 error annotations. If there are more than that, use only the most severe ones.
* Make the final annotation set as concise as possible in terms of number of error annotations.
Dont use any markdown formatting. Generate merged error annotations in this format, without any additional text:
Error 1:
Location: <span of text from the output, or None if not applicable>
Explanation: <explanation>
Severity: <severity level>
### Model output
{{ output }}
### Error annotations
{{ annotations }}
Listing 2: Supervisor prompt template for merging of annotation
### Task
Your task is to evaluate a model output for {{ task_name }} task with respect to {{ aspect_name }}. {{ extra_task_info }}
### Aspect Definition
{{ aspect_name }} - {{ aspect_definition }}
### Dialogue history
{{ input }}
{if context %}
### Knowledge
{{ context }}
{endif %}
### Response
{{ output }}
### Instructions
For any error in the output, identify its location, assign a severity level and provide an explanation. Report at most 8 errors. If there are more errors, report only the most severe ones. Finally, provide an overall score between 0 and 100 for {{ aspect_name }} of the output.
Listing 3: Prompt template for the fine-tuned LLM
Metric Consistency Coherence Relevance Fluency Average
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
ROUGE-1 0.167 0.126 0.160 0.130 0.115 0.094 0.326 0.252 0.192 0.150
ROUGE-2 0.184 0.139 0.187 0.155 0.159 0.128 0.290 0.219 0.205 0.161
ROUGE-L 0.128 0.099 0.115 0.092 0.105 0.084 0.311 0.237 0.165 0.128
BERTScore 0.151 0.122 0.285 0.220 0.302 0.232 0.186 0.154 0.231 0.182
MOVERSscore 0.159 0.118 0.157 0.127 0.129 0.105 0.318 0.244 0.191 0.148
BARTScore 0.266 0.220 0.474 0.367 0.318 0.243 0.258 0.214 0.329 0.261
UniEval 0.446 0.371 0.575 0.442 0.426 0.325 0.449 0.371 0.474 0.377
G-Eval (GPT-3.5) 0.386 0.318 0.440 0.335 0.385 0.293 0.424 0.347 0.409 0.323
G-Eval (GPT-4) 0.507 0.425 0.582 0.457 0.548 0.433 0.455 0.378 0.523 0.423
GPT-3.5 0.393 0.331 0.459 0.371 0.455 0.363 0.355 0.296 0.415 0.340
GPT-4 0.531 0.464 0.540 0.434 0.491 0.395 0.480 0.409 0.511 0.426
Llama 3.1 8B 0.181 0.165 0.176 0.150 0.156 0.127 0.232 0.218 0.186 0.165
X-Eval 0.428 0.340 0.530 0.382 0.500 0.361 0.461 0.365 0.480 0.362
Prometheus 0.150 0.137 0.150 0.126 0.164 0.138 0.189 0.168 0.163 0.142
AUTO-J 0.131 0.121 0.245 0.203 0.262 0.222 0.154 0.141 0.198 0.172
TIGERScore 0.427 0.387 0.381 0.318 0.366 0.304 0.363 0.327 0.384 0.334
InstructScore 0.232 0.213 0.328 0.276 0.211 0.179 0.260 0.237 0.258 0.226
Themis 0.600 0.566 0.566 0.485 0.474 0.412 0.571 0.533 0.553 0.499
OpeNLGaugeens 0.548 0.470 0.604 0.462 0.513 0.389 0.470 0.389 0.534 0.427
- Command R+ 104B 0.633 0.603 0.239 0.203 0.360 0.302 0.347 0.320 0.395 0.357
- Gemma 2 27B 0.459 0.421 0.455 0.386 0.428 0.358 0.427 0.386 0.442 0.388
- Llama 3.1 Nemotron 70B 0.559 0.517 0.469 0.391 0.419 0.356 0.358 0.325 0.451 0.397
- Mistral Large 2 123B 0.627 0.590 0.528 0.434 0.456 0.375 0.398 0.359 0.502 0.439
- Qwen 2.5 70B 0.567 0.521 0.525 0.433 0.388 0.317 0.433 0.389 0.478 0.415
OpeNLGaugeft 0.527 0.453 0.561 0.441 0.514 0.408 0.407 0.349 0.502 0.413
Table 12: Segment-level Spearman ($\rho$) and Kendall ($\tau$) correlations of different metrics on SummEval.
Metric Groundedness Coherence Engagingness Naturalness Average
$r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$
ROUGE-L 0.193 0.203 0.176 0.146 0.295 0.300 0.310 0.327 0.243 0.244
BLEU-4 0.131 0.235 0.180 0.175 0.232 0.316 0.213 0.310 0.189 0.259
METEOR 0.250 0.302 0.212 0.191 0.367 0.439 0.333 0.391 0.290 0.331
BERTScore 0.214 0.233 0.226 0.209 0.317 0.335 0.291 0.317 0.262 0.273
USR 0.416 0.377 0.337 0.325 0.456 0.465 0.222 0.447 0.358 0.403
UniEval 0.602 0.455 0.455 0.330 0.573 0.430 0.577 0.453 0.552 0.417
G-Eval (GPT-3.5) 0.586 0.567 0.519 0.544 0.660 0.691 0.532 0.539 0.574 0.585
G-Eval (GPT-4) 0.531 0.551 0.594 0.605 0.627 0.631 0.549 0.565 0.575 0.588
GPT-3.5 0.653 0.581 0.550 0.531 0.651 0.648 0.515 0.550 0.592 0.578
GPT-4 0.810 0.786 0.680 0.680 0.822 0.779 0.769 0.739 0.770 0.746
Llama 3.1 8B 0.374 0.362 0.225 0.228 0.415 0.408 0.222 0.238 0.309 0.309
X-Eval 0.734 0.728 0.558 0.622 0.449 0.593 0.417 0.478 0.539 0.605
Prometheus 0.437 0.412 0.451 0.465 0.495 0.473 0.355 0.384 0.435 0.434
AUTO-J 0.339 0.357 0.452 0.449 0.490 0.459 0.425 0.437 0.427 0.425
TIGERScore 0.137 0.138 0.417 0.438 0.328 0.333 0.455 0.477 0.334 0.346
InstructScore 0.140 0.102 0.299 0.297 0.264 0.233 0.374 0.332 0.269 0.241
Themis 0.778 0.761 0.639 0.644 0.790 0.766 0.727 0.729 0.733 0.725
OpeNLGaugeens 0.704 0.697 0.622 0.621 0.675 0.692 0.599 0.604 0.649 0.653
- Command R+ 104B 0.383 0.368 0.463 0.453 0.262 0.259 0.421 0.398 0.386 0.374
- Gemma 2 27B 0.332 0.366 0.465 0.481 0.549 0.562 0.489 0.515 0.459 0.484
- Llama 3.1 Nemotron 70B 0.781 0.791 0.600 0.630 0.655 0.686 0.506 0.532 0.621 0.645
- Mistral Large 2 123B 0.658 0.648 0.541 0.554 0.645 0.659 0.509 0.529 0.586 0.596
- Qwen 2.5 70B 0.467 0.460 0.480 0.521 0.500 0.516 0.468 0.488 0.496 0.514
OpeNLGaugeft 0.485 0.538 0.531 0.575 0.622 0.650 0.522 0.547 0.540 0.578
Table 13: Segment-level Pearson ($r$) and Spearman ($\rho$) correlations of different metrics on TopicalChat.
Metric SFRES SFHOT Average
Inf. Nat. Inf. Nat.
ROUGE-1 0.115 0.170 0.118 0.196 0.150
ROUGE-L 0.103 0.169 0.110 0.186 0.142
BERTScore 0.156 0.219 0.135 0.178 0.172
MOVERScore 0.153 0.190 0.172 0.242 0.189
BARTScore 0.238 0.289 0.235 0.288 0.263
UniEval 0.225 0.333 0.249 0.320 0.282
GPTScore 0.232 0.190 0.184 0.036 0.161
G-Eval 0.189 0.351 0.198 0.338 0.269
GPT-3.5 0.304 0.385 0.306 0.242 0.294
GPT-4 0.213 0.405 0.320 0.302 0.359
Llama 3.1 8B 0.003 0.081 0.071 0.094 0.108
Prometheus 0.161 0.150 0.173 0.169 0.211
Auto-J 0.179 0.084 0.141 0.176 0.127
TIGERScore 0.160 0.221 0.200 0.215 0.204
InstructScore 0.194 0.300 0.247 0.222 0.273
Themis 0.298 0.395 0.333 0.259 0.380
OpeNLGaugeens 0.234 0.415 0.205 0.341 0.299
- Command R+ 104B 0.236 0.299 0.047 0.210 0.198
- Gemma 2 27B 0.254 0.334 0.275 0.317 0.295
- Llama 3.1 Nemotron 70B 0.178 0.284 0.226 0.281 0.242
- Mistral Large 2 123B 0.129 0.359 0.215 0.311 0.254
- Qwen 2.5 70B 0.226 0.347 0.245 0.315 0.283
OpeNLGaugeft 0.354 0.354 0.238 0.315 0.315
Table 14: Segment-level Spearman ($r$) correlations of different metrics on SFRES and SFHOT. Inf. = informativeness, Nat. = naturalness.
Metric Coh. Rel. Eng. Emp. Sur. Com. Avg.
BLEU 0.539 0.514 0.483 0.410 0.471 0.516 0.489
ROUGE-1 0.567 0.518 0.529 0.450 0.490 0.591 0.524
METEOR 0.560 0.522 0.510 0.435 0.488 0.555 0.512
MoverScore 0.551 0.523 0.495 0.418 0.478 0.530 0.499
BERTScore 0.566 0.531 0.520 0.441 0.488 0.563 0.518
BARTScore 0.501 0.467 0.465 0.416 0.436 0.488 0.462
Llama 3.1 8B 0.119 0.344 0.154 0.094 0.080 0.212 0.167
OpeNLGaugeens 0.528 0.559 0.538 0.434 0.343 0.591 0.499
- Command R+ 104B 0.412 0.383 0.420 0.330 0.227 0.344 0.353
- Gemma 2 27B 0.453 0.445 0.461 0.356 0.323 0.521 0.427
- Llama 3.1 Nemotron 70B 0.500 0.508 0.497 0.380 0.332 0.565 0.464
- Mistral Large 2 123B 0.372 0.490 0.425 0.373 0.334 0.507 0.417
- Qwen 2.5 72B 0.407 0.484 0.427 0.277 -0.012 0.507 0.348
OpeNLGaugeft 0.448 0.523 0.517 0.444 0.404 0.555 0.482
Table 15: Segment-level Pearson ($r$) correlations of various metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
Metric Coh. Rel. Eng. Emp. Sur. Com. Avg.
BLEU 0.339 0.292 0.356 0.315 0.299 0.414 0.336
ROUGE-1 0.389 0.330 0.416 0.354 0.355 0.503 0.391
METEOR 0.378 0.310 0.412 0.366 0.354 0.505 0.387
MoverScore 0.392 0.385 0.420 0.331 0.321 0.473 0.387
BERTScore 0.372 0.355 0.415 0.356 0.320 0.469 0.381
BARTScore 0.259 0.249 0.291 0.287 0.227 0.294 0.268
Llama 3.1 8B 0.093 0.334 0.140 0.086 0.064 0.184 0.150
OpeNLGaugeens 0.393 0.474 0.452 0.367 0.276 0.489 0.409
- Command R+ 104B 0.325 0.362 0.372 0.320 0.215 0.348 0.324
- Gemma 2 27B 0.381 0.383 0.400 0.310 0.278 0.445 0.366
- Llama 3.1 Nemotron 70B 0.429 0.445 0.427 0.328 0.263 0.464 0.393
- Mistral Large 2 123B 0.294 0.415 0.389 0.349 0.337 0.474 0.376
- Qwen 2.5 72B 0.344 0.423 0.364 0.262 -0.006 0.416 0.301
OpeNLGaugeft 0.416 0.498 0.464 0.391 0.319 0.460 0.425
Table 16: Segment-level Spearman correlations ($\rho$) of various metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
Metric Coh. Rel. Eng. Emp. Sur. Com. Avg.
BLEU 0.248 0.209 0.260 0.230 0.220 0.305 0.245
ROUGE-1 0.287 0.237 0.306 0.260 0.262 0.376 0.288
METEOR 0.278 0.224 0.303 0.269 0.261 0.377 0.285
MoverScore 0.289 0.280 0.308 0.242 0.236 0.353 0.285
BERTScore 0.273 0.257 0.304 0.260 0.234 0.348 0.279
BARTScore 0.185 0.177 0.209 0.206 0.164 0.212 0.192
Llama 3.1 8B 0.079 0.285 0.116 0.073 0.055 0.152 0.127
OpeNLGaugeens 0.307 0.367 0.350 0.280 0.208 0.378 0.315
- Command R+ 104B 0.270 0.297 0.300 0.253 0.171 0.277 0.261
- Gemma 2 27B 0.329 0.328 0.343 0.267 0.241 0.384 0.315
- Llama 3.1 Nemotron 70B 0.369 0.377 0.364 0.272 0.221 0.388 0.332
- Mistral Large 2 123B 0.253 0.354 0.334 0.301 0.284 0.405 0.322
- Qwen 2.5 72B 0.294 0.360 0.307 0.222 -0.003 0.353 0.255
OpeNLGaugeft 0.341 0.400 0.377 0.323 0.257 0.377 0.346
Table 17: Segment-level Kendall ($\tau$) correlations of various metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
Metric Fluency Meaning Simplicity Average
BLEU 0.460 0.622 0.438 0.507
SARI 0.335 0.534 0.366 0.412
BERTScore 0.636 0.682 0.614 0.644
LENS 0.816 0.662 0.733 0.737
Llama 3.1 8B 0.373 0.528 0.313 0.405
OpeNLGaugeens 0.840 0.864 0.770 0.825
- Command R+ 104B 0.704 0.787 0.601 0.697
- Gemma 2 27B 0.755 0.769 0.688 0.737
- Llama 3.1 Nemotron 70B 0.778 0.822 0.660 0.753
- Mistral Large 2 123B 0.705 0.744 0.735 0.728
- Qwen 2.5 72B 0.771 0.829 0.730 0.776
OpeNLGaugeft 0.801 0.851 0.716 0.789
Table 18: Segment-level Pearson ($r$) correlations of different metrics on Wiki-DA.
Figure 4: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the Wiki-DA dataset. Specific model combinations are represented by the colored patches.
Figure 5: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the QAGS dataset.
Figure 6: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the SummEval dataset.
Figure 7: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the HANNA dataset.
Figure 8: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the TopicalChat dataset.
Figure 9: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the SFRES/SFHOT dataset.
Figure 10: Results for Ablation 1 on QAGS and TopicalChat. The LLMs are instructed to provide both integer overall scores (1–5) and integer severity levels (1–5). The plotted values represent differences in Spearman’s $\rho$ correlations with human scores between the original prompt and the ablation. For TopicalChat, Coh. = coherence, Eng. = engagingness, Gro. = groundedness, Nat. = naturalness, Avg. = average of all aspects.
Figure 11: Results for Ablation 2 on QAGS and TopicalChat. The LLMs are instructed to provide integer overall scores (1–5), and categorical severity levels on the following scale: Neutral, Minor, Moderate, Major, Critical.
Figure 12: Results for Ablation 3 on QAGS and TopicalChat. The LLMs are instructed to provide categorical overall scores on the scale described in Section 4, and categorical severity levels on the following scale: Neutral, Minor, Moderate, Major, Critical.
Figure 13: Results for Ablation 4 on QAGS and TopicalChat, where the rules section is removed from the prompt.
Data: Adolfo_Suárez_Madrid–Barajas_Airport | operatingOrganisation | ENAIRE
Adolfo_Suárez_Madrid–Barajas_Airport | runwayLength | 3500.0
Adolfo_Suárez_Madrid–Barajas_Airport | elevationAboveTheSeaLevel | 610.0
Adolfo_Suárez_Madrid–Barajas_Airport | runwayName | "14L/32R"
Text: Adolfo Suárez Madrid–Barajas Airport is a major international airport located near the heart of the city, offering a wide range of air services.
Evaluation aspect: Faithfulness – Extent to which the information in the text is supported by thedata Explanation: This phrase is considered an error because the provided data does not mention the airport’s size (major/international), location relative to the city (near the heart), or any information that could imply these details. The text introduces unsupported information, negatively affecting faithfulness. Severity: 4 Explanation: Similar to Error 1, this phrase introduces information not present in the provided data. There is no mention of the types or variety of air services offered by the airport, making this an unsupported claim that affects faithfulness. Severity: 3 Overall Score: Poor (2/5) Explanation of the Score: The generated text contains significant errors related to faithfulness, as it introduces substantial information not supported by the provided data. While the text does not contradict any data points (since it doesn’t directly address them), its lack of adherence to the given facts and introduction of unsubstantiated claims warrant a ‘Poor’ rating. For an ‘Excellent’ score, the text should have strictly adhered to the provided data, which includes specific details about the airport’s operating organisation, runway length, elevation, and runway name, none of which were accurately represented or utilized in the generated text.
Figure 14: Example error span annotation provided by OpeNLGauge for the RDF-to-text task. The text generated from the data is evaluated for faithfulness.
Emotional Context: A is terrified Dialogue History: A: I moved into this new house and was woken up with this horrible screeching noise in the middle of the night. B: Terrifying! What was it? A: It turned out to be the sump pump, I guess it was old. But yes, it scared the daylights out of me! B: Oh good. I bet! I wouldn’t have been able to go back to sleep after that! Do you have to replace the pump? A: I did, yeah. It was actually still in working condition but I couldn’t stand the sound and I’m a light sleeper. No more scary noises at 3am! Response: that ’s great! i’m glad you’re able to deal with it now! Evaluation aspect: Empathy – Extent to which the response shows understanding of the feelings of the person talking about their experience Explanation: The response acknowledges the resolution of the issue but fails to explicitly acknowledge A’s emotional distress (being terrified) caused by the event. Using a casual, positive phrase like "that’s great" might come across as dismissive of the initial fear experienced. Severity: 3 Explanation: While this part of the response shows some form of relief for A’s current situation, it doesn’t directly address or validate A’s past emotional state (terror) related to the incident. It focuses more on the practical outcome rather than offering empathy for the emotional experience. Severity: 2 Overall Score: Fair (3/5) Explanation of the Score: The response demonstrates a basic understanding that the issue is resolved and expresses relief for the person, which shows some level of consideration. However, it lacks direct acknowledgment and validation of A’s terrified emotional state, preventing it from being scored higher in empathy. The errors identified, while present, do not completely negate the empathetic intent, hence a ’Fair’ score rather than ’Poor’ or ’Unacceptable’.
Figure 15: Example error span annotation provided by OpeNLGauge for the dialogue response generation task. The response is evaluated for empathy. In addition to dialogue history, the input also includes the emotional context.
Summary: By A new study that college are more likely to engage in energy - binge drinking when alcohol than non - . Evaluation aspect: Fluency – Formal quality of individual sentences of the summary. A fluent sen-tence should be grammatical, natural and easy to understand. Explanation: The sentence is fragmented and lacks proper structure. It starts with an unnecessary preposition ("By") and contains several grammatical errors, such as missing words and incorrect phrasing. This makes the sentence difficult to understand and unnatural. Severity: 5 Overall Score: Unacceptable (1/5) Explanation of the Score: The summary is very disfluent due to significant grammatical errors and lack of coherence, making it extremely difficult to comprehend.
Figure 16: Example error span annotation provided by OpeNLGauge for the summarization task. The summary is evaluated for fluency. Note that source text is not included as the input for fluency evaluation.