### v3

# OpeNLGauge: An Explainable Metric for NLG Evaluation with Open-Weights LLMs

Abstract
Large Language Models (LLMs) have demonstrated great potential as evaluators of NLG systems, allowing for high-quality, reference-free, and multi-aspect assessments. However, existing LLM-based metrics suffer from two major drawbacks: reliance on proprietary models to generate training data or perform evaluations, and a lack of fine-grained, explanatory feedback. We introduce OpeNLGauge, a fully open-source, reference-free NLG evaluation metric that provides accurate explanations based on individual error spans. OpeNLGauge is available as a two-stage ensemble of larger open-weight LLMs, or as a small fine-tuned evaluation model, with confirmed generalizability to unseen tasks, domains and aspects. Our extensive meta-evaluation shows that OpeNLGauge achieves competitive correlation with human judgments, outperforming state-of-the-art models on certain tasks while maintaining full reproducibility and providing explanations more than twice as accurate.
OpeNLGauge: An Explainable Metric for NLG Evaluation with Open-Weights LLMs
Ivan Kartáč  and  Mateusz Lango  and  Ondřej Dušek Charles University, Faculty of Mathematics and Physics Institute of Formal and Applied Linguistics Prague, Czechia {kartac,lango,odusek}@ufal.mff.cuni.cz
1 Introduction
Evaluating Natural Language Generation (NLG) systems remains a challenging research problem. Traditional overlap-based metrics, such as BLEU papineni2002bleu and ROUGE lin-2004-rouge, are still widely used but exhibit limited correlation with human judgments, particularly when assessing modern NLG systems novikova-etal-2017-need. With the rise of pre-trained language models, the research community began to shift toward model-based metrics that better capture semantic similarity, yet their performance was still unsatisfactory yan-etal-2023-bleurt; glushkova-etal-2023-bleu.
Figure 1: The ensemble metric OpeNLGaugeens and its distilled version OpeNLGaugeft.
Recently, Large Language Models (LLMs) have demonstrated remarkable potential in imitating human evaluation of generated text jiang2023tigerscore; xu-etal-2023-instructscore; hu2024themis. LLM-based metrics are often general enough to evaluate diverse NLG tasks and can provide high evaluation performance without the need for reference texts liu-etal-2023-g. They also offer better adaptability to evaluate specific aspects of generated text and are able to evaluate beyond semantic correctness, taking into account aspects such as factual consistency or relevance gu2025surveyllmasajudge. However, existing LLM-based NLG metrics lack two important features: (1) they are not fully open-source, with most relying on data generated by proprietary LLMs such as GPT-4 openai2024gpt4technicalreport, and (2) they do not provide precise explanations for their evaluations, with most of them producing only overall scores, or explanations limited to a single short comment (see Table 1 for details).
In this paper, we introduce OpeNLGauge – a versatile, reference-free metric for NLG tasks that provides precise, error-span-based explanations (Figure 1). Unlike existing LLM-based metrics, OpeNLGauge is built entirely on open-weight models and does not rely on human-annotated datasets. It supports a wide range of NLG tasks, including data-to-text, summarization or story generation. Moreover, it allows for fine-grained evaluation of specific, customizable aspects of the generated text such as faithfulness, fluency or coherence. It can also evaluate other user-defined aspects appropriate for a given application.
Our contributions are as follows:
  • We introduce an effective prompting strategy and a two-stage LLM ensemble to obtain high-quality evaluations of NLG outputs. The proposed OpeNLGaugeens demonstrates higher correlations with human judgments on some NLG tasks than previously proposed metrics based on proprietary LLMs.
  • We collect outputs from over 100 NLG systems on 15 NLG tasks and use OpeNLGaugeens to annotate them to construct a comprehensive dataset for training open NLG metrics.
  • The collected dataset is used to construct OpeNLGaugeft – a fine-tuned version of the smallest model from the Llama 3.1 family grattafiori2024llama3herdmodels, which is able to provide accurate explanations and reliable quality assessments of NLG outputs in a more cost-effective way.
  • We perform an extensive meta-evaluation of the proposed metrics on seven datasets covering five NLG tasks. The experimental analysis includes human evaluation of explanation quality, comparison of different score aggregation methods, ablation experiments, and evaluation on tasks, domains, and aspects not seen during training.
Our experiments show that OpeNLGauge achieves higher explanation quality than other metrics trained on data generated by proprietary LLMs, while also delivering strong evaluation performance. On tasks such as summarization, OpeNLGauge achieves higher correlations with human judgments than metrics based on the strong proprietary GPT-4 model. While the state-of-the-art metric, Themis – which leverages both human-annotated data and data from proprietary LLMs – remains superior in terms of correlations with human judgments, OpeNLGauge proves highly competitive and even outperforms it on certain tasks, while also providing more detailed error explanations.
All our experiments were conducted using only quantized open-weight models and two GPUs with 48GB of VRAM each, ensuring that the results can be reproduced in many AI research labs. Our code and data are available on Github. [^1] https://github.com/ivankartac/OpeNLGauge
2 Related Work
Metric No Ref. Aspects Open Err. Span
G-Eval
Prometheus
Auto-J
InstructScore
TIGERScore
Themis
[0.5pt/2pt] OpeNLGauge
Table 1: Comparison of properties in different LLM-based metrics for NLG: reference-free (No Ref.), customizable aspects (Aspects), built exclusively using open-weight LLMs (Open), and producing precise explanations with error-span annotation (Err. Span). The metrics compared are: G-Eval liu-etal-2023-g, Prometheus (kim2023prometheus), Auto-J (li2023generative), InstructScore (xu-etal-2023-instructscore), TIGERScore (jiang2023tigerscore), Themis (hu2024themis) and OpeNLGauge (this work).
Context: [SUMMARY OF THE STORY] Question: What is Dr. Heywood Floyd’s mission on the Clavius Base? Answer: According to the summary provided[1], Dr. Heywood Floyd’s mission on the Clavius Base is to investigate a recently found artifact buried four million years ago. Specifically, his mission is to ride in a Moonbus to the artifact, which is a monolith identical to the one encountered by the man-apes in the summary.[2] Evaluation aspect:Conciseness – Extent to which the answer is concise and to the point [1] Explanation: This introductory phrase is an error affecting conciseness as it adds unnecessary words to the answer. The response could directly state Floyd’s mission without referencing the source (summary), given the context of the question. Severity: 2 [2] Explanation: This phrase is considered an error affecting conciseness because it provides additional, unnecessary details about the mission. The initial sentence already clearly states Floyd’s mission, making the subsequent elaboration redundant and wordy. Severity: 3 Overall score: Fair (3/5) Explanation of the score: While the generated answer generally addresses the question, its conciseness is compromised by the inclusion of redundant information and an unnecessary introductory phrase. Removing these elements would enhance the answer’s directness and efficiency, potentially elevating the score to ‘Good’.
Figure 2: Example error span annotation provided by OpeNLGauge for the narrative question answering task. The answer to the question, grounded in the story summary, is evaluated for conciseness.
Although NLG has traditionally been evaluated using simple word-overlap-based metrics such as BLEU, these are known to have low correlations with human judgments novikova-etal-2017-need; reiter2018structured. This improved somewhat with the use of trained models for metrics in the past few years yuan2021bartscore; zhong2022towards; mehri2020usr, but the correlations remained moderate. Recently, numerous studies explored the application of LLMs in NLG evaluation. A prominent line of research focuses on leveraging proprietary LLMs such as GPT-4 openai2024gpt4technicalreport, with direct prompting for an overall score fu2024gptscore; kocmi2023large or even annotating error spans with categories, which has been explored in machine translation kocmi2023gemba; fernandes2023devil; lu2023error. However, using proprietary models is costly and comes with a reproducibility penalty, as some LLM versions become unavailable or are modified in a non-transparent way (Chen2024How). Another concern is data leakage, which affects results, but cannot be controlled in the case of proprietary models balloccu2024leak.
Many recent LLM-based metrics, such as Themis (hu2024themis), TIGERScore (jiang2023tigerscore), InstructScore xu-etal-2023-instructscore or Prometheus kim2023prometheus, are built on open LLMs, but in fact they all rely on proprietary LLMs such as GPT-4 openai2024gpt4technicalreport to generate, filter, or annotate their training data. Therefore, some of these metrics can be viewed as sophisticated knowledge distillation methods from proprietary to open-weight LLMs. This retains the reproducibility disadvantage, as reconstructing the metric from scratch or adapting it to a new task requires access to closed-source LLMs.
While several metrics based on open-weight LLMs provide some level of interpretability, this is often limited to a short, free-text review of the evaluated output hu2024themis; kim2023prometheus. Fine-grained error-span annotation offers several advantages over singular scores and comments: (1) it can be easily processed automatically, allowing its use in post-processing steps or to provide feedback to the model or training algorithms, (2) it offers greater precision and clarity, as errors are associated with a particular part of the output, making it easier to find and correct issues, (3) it is more human-like, resembling the output of human annotation schemes such as MQM (10.1162/tacl_a_00437) or ESA (kocmi2024error). The alignment with human evaluation allows easier comparison to humans or even use as pre-annotation, helping to accelerate human annotation process. While some metrics do provide annotation on the error level, these tend to use closed LLMs and their scope is generally limited to single NLG tasks, such as the machine translation approaches mentioned above or kasner2024beyond’s work on data-to-text generation.
Scoring outputs with a single LLM may introduce bias (zheng2023judging; panickssery2024llm), which can be alleviated by combining multiple LLMs as evaluators. verga2024replacingjudgesjuriesevaluating apply an ensemble of mostly proprietary LLMs, aggregating their scores either by majority voting over binary ratings, or by averaging for ordinal scores. For pairwise evaluation, li2024prdpeerrankdiscussion propose two methods which combine preferences of multiple LLMs, including an iterative multi-agent discussion.
3 Problem Statement
We formulate the problem of evaluating the output of an NLG system while providing error-based explanations as follows. Given an input $x$, an output $y$, and an evaluation aspect $a$, the task is to return a tuple $\langle z,\{e_{1},...,e_{n}\}\rangle$, where $z$ is a numeric score assigned to $y$, and $\{e_{1},...,e_{n}\}$ represents a set of error annotations. Each error annotation $e_{i}=\langle s_{i},t_{i},l_{i}\rangle$ includes a span of text $s_{i}\in y$ corresponding to the problematic segment, a textual explanation $t_{i}$, and a severity level $l_{i}$. Although the term error is used throughout this work, it should be understood in a broader sense as any issue in the text related to the evaluated aspect $a$. Examples of outputs provided by OpeNLGauge are presented in Figure 2 and in Appendix L.
4 Open LLM Ensemble as Evaluator
To achieve a high-quality evaluation of NLG outputs, we propose OpeNLGaugeens, a two-stage ensemble of open-weight LLMs. The ensemble consists of $n$ annotator models, which perform independent analyses of the provided NLG output, and a consolidator model, which is responsible for merging their results and filtering inaccuracies. A high-level overview of the approach is presented in Figure 1.
Although our approach requires multiple LLMs, using the ensemble with a handful of models ($n=5$ in our experiments) is still less computationally demanding than sampling multiple outputs to obtain statistical estimates, as required by some metrics, e.g. G-Eval ($n=20$, liu-etal-2023-g). Furthermore, we only use quantized LLMs to limit the computational requirements.
Annotator models
The annotator models are open-weight LLMs prompted to identify error spans in the text and to provide detailed explanations and severity levels for each error. This approach facilitates the interpretability of the evaluation process, but also acts as a chain-of-thought mechanism 10.5555/3600270.3602070, helping the model to ground its decisions in a structured reasoning path.
The full annotator model prompt is provided in Appendix F. It contains the description of the evaluated task (e.g., data-to-text), the definition of the evaluated aspect, and a template for the model’s response. Since LLMs are known to confuse different evaluation aspects (hu-etal-2024-llm), the prompt also contains several rules that instruct the model to remain focused on the specific evaluation aspect, not to make additional assumptions, and to justify any score lower than the maximum by at least one identified error. Inspired by liu-etal-2023-g, we also include detailed steps for error identification.
Furthermore, we provide a description of the overall scoring scale, including an explanation of the lowest and highest scores. The scale is presented as categorical (Unacceptable < Poor < Fair < Good < Excellent), based on the intuition that adjectival categorical scales may be easier for language models to interpret than numerical scales. We use an integer severity scale (1-5) for scoring individual errors in order to avoid confusion with the overall scoring scale. [^2] A categorical scale was also initially used for error severity. However, we found that LLMs sometimes confused the error severity scale with the overall score scale.
Finally, the prompt contains the input that was originally used to generate the evaluated output (e.g., the source text for summarization or the input data in data-to-text). Some tasks may involve multiple inputs; for instance, evaluating knowledge-grounded question answering requires knowledge of both the question and the context. In such cases, the context is also presented to the model under separate headers. Although structured formats like JSON might make parsing of the output easier, prompting LLMs to reason within strict structured outputs has been shown to impair their performance (tam-etal-2024-speak; kellner2024guiding). Therefore, the models are instructed to produce textual outputs.
Consolidator model
The final score of the NLG output is computed as a simple average of the scores provided by the annotator models. However, to meet the requirement of explainable output, the error analyses of multiple annotators need to be unified. This is the task of the consolidator LLM.
This open-weight LLM is instructed to: (1) merge errors detected by multiple models, (2) unify their output format, and (3) clean up the annotations. Specifically, the model is instructed to merge all error annotations that refer to the same issue at approximately the same location in the text, while maintaining annotation granularity. It is also prompted to fix potential deviations from the expected output format. To simplify the cleaning process, error analyses produced by outlier annotator models [^3] The annotation is considered to be an outlier if the score provided by the annotator differs from the mean score by at least two standard deviations, and this difference is at least 1. are filtered out from the input to the consolidator model. The full prompt for the consolidator model is provided in Appendix F.
5 Training a Distilled Model
5.1 Synthetic data generation
To distill knowledge from the ensemble, we collected the outputs of over a hundred NLG systems and applied OpeNLGaugeens to produce synthetic evaluation data. We include NLG outputs produced on 15 datasets covering five task categories and almost 40 aspects. Table 2 shows the basic statistics of the constructed dataset.
Task Src. Sys. Asp. Examples
Summarization 5 39 6 12,070
Data-to-text 4 29 7 07,894
Dialogue 3 36 9 10,074
Story Generation 1 9 6 03,200
Question Answering 2 15 11 04,849
Table 2: Training dataset statistics for OpeNLGaugeft: number of source datasets (Src.), systems (Sys.), evaluation aspects (Asp.) and training examples for each task.
The NLG outputs included cover the following tasks: data-to-text, summarization, question answering, dialogue response generation and story generation. These were chosen to represent a diverse range of tasks, each with unique objectives, input-output relationships and evaluation aspects. Note that the underlying datasets are used only as inputs for NLG systems and do not include human evaluations of any kind.
For each task, we select a set of relevant evaluation aspects with their definitions. We consider two aspects distinct if they are associated with different tasks. For instance, coherence in dialogue refers to coherence of the response with respect to the dialogue history, while in summarization it refers to the internal coherence of a summary. We list the datasets and aspects for each task in Appendix A.
To obtain output texts with varying quality and diverse types of errors, we sample outputs from a variety of systems, ranging from rule-based approaches to state-of-the-art LLMs. For older systems, we use pre-generated outputs from existing datasets, while outputs from more recent systems including LLMs are newly generated. An overview of the evaluated systems is shown in Appendix B.
To limit computational requirements for dataset generation, we apply a sampling procedure that ensures data diversity and broad coverage of different NLG systems and aspects while significantly reducing dataset size. For each input, we randomly sample the outputs of $N$ systems, followed by sampling $M$ aspects for each input-output pair. The sampling includes all aspects described in Appendix A and all systems listed in Appendix B. This results in $N\,\times\,M$ (input, output, aspect) triples for each input, which are then passed to OpeNLGaugeens to obtain synthetic annotations. For most tasks, we set $N=4$ and $M$ = 3. This sampling strategy aims for a balanced distribution of inputs, NLG system outputs and evaluation aspects. It also ensures exposure to different outputs for the same input, i.e., it should not prime models to evaluate based solely on patterns in the inputs. Finally, by presenting multiple aspects for the same input-output pair, models are encouraged to learn differences in output quality between different evaluation aspects.
To keep the merged evaluation outputs internally consistent, we remove outliers before merging, as described in Section 4. Table 12 in Appendix D shows the proportions of outliers detected for all LLMs and task categories (3.4% on average).
5.2 Fine-tuning Procedure
We use the dataset described in Section 5.1 for supervised fine-tuning of a specialized LLM evaluator, with an instruction-tuned version of Llama 3.1 8B as the backbone. To avoid training the LLM to predict floating-point scores, we convert them to integers in the range 0-100 and then bin them to the nearest multiple of five. This extends the output space from five to twenty values (instead of 100), which is a trade-off between greater granularity in predictions and manageable task complexity.
As the model is expected to learn the task from training data, we used a simpler prompt template for fine-tuning, with a brief description of the task, the definition of the evaluated aspect, and the input and output to be evaluated (see Appendix F).
Metric QAGS-CNN/DM QAGS-XSUM Average
$r$ $\rho$ $\tau$ $r$ $\rho$ $\tau$ $r$ $\rho$ $\tau$
ROUGE-1 0.338 0.318 0.248 -0.008 -0.049 -0.040 0.165 0.134 0.104
ROUGE-2 0.459 0.418 0.333 0.097 0.083 0.068 0.278 0.250 0.200
ROUGE-L 0.357 0.324 0.254 0.024 -0.011 -0.009 0.190 0.156 0.122
BERTScore 0.576 0.505 0.399 0.024 0.008 0.006 0.300 0.256 0.202
MoverScore 0.414 0.347 0.271 0.054 0.044 0.036 0.234 0.195 0.153
FactCC 0.416 0.484 0.376 0.297 0.259 0.212 0.356 0.371 0.294
QAGS 0.545 - - 0.175 - - 0.375 - -
BARTScore 0.732 0.680 0.555 0.175 0.171 0.139 0.454 0.425 0.347
UniEval 0.682 0.662 0.532 0.461 0.488 0.399 0.572 0.575 0.466
G-Eval (GPT-3.5) 0.477 0.516 0.410 0.211 0.406 0.343 0.344 0.461 0.377
G-Eval (GPT-4) 0.631 0.685 0.591 0.558 0.537 0.472 0.595 0.611 0.532
LLM Evaluation (GPT-3.5) 0.454 0.514 0.417 0.279 0.348 0.295 0.366 0.431 0.356
LLM Evaluation (GPT-4) 0.735 0.746 0.626 0.541 0.528 0.439 0.638 0.637 0.532
Auto-J 0.291 0.238 0.214 0.225 0.214 0.203 0.258 0.226 0.209
TIGERScore 0.574 0.562 0.479 0.424 0.445 0.412 0.499 0.504 0.446
InstructScore 0.287 0.278 0.233 -0.096 -0.134 -0.119 0.095 0.072 0.057
Themis 0.747 0.761 0.680 0.599 0.607 0.546 0.673 0.684 0.613
OpeNLGaugeens 0.738 0.753 0.627 0.630 0.624 0.531 0.684 0.689 0.579
• Command R+ 104B 0.676 0.675 0.617 0.540 0.541 0.515 0.608 0.608 0.566
• Gemma 2 27B 0.579 0.646 0.579 0.592 0.614 0.563 0.585 0.630 0.571
• Llama 3.1 Nemotron 70B 0.705 0.733 0.650 0.587 0.586 0.540 0.646 0.659 0.595
• Mistral Large 2 123B 0.658 0.704 0.635 0.577 0.570 0.541 0.617 0.637 0.588
• Qwen 2.5 72B 0.678 0.720 0.635 0.568 0.569 0.526 0.623 0.644 0.581
Llama 3.1 8B 0.275 0.242 0.219 0.218 0.230 0.218 0.247 0.236 0.219
OpeNLGaugeft 0.668 0.695 0.584 0.607 0.607 0.524 0.638 0.651 0.554
Table 3: Segment-level Pearson ($r$), Spearman ($\rho$) and Kendall ($\tau$) correlations of different metrics for factual consistency on QAGS. The best correlations are highlighted in bold, the second best are underlined.
6 Experimental Setup
6.1 Ensemble
The ensemble consists of six open-weight LLMs: five annotators and one consolidator. Each selected LLM is distributed under a license that permits at least non-commercial use and allows the model’s outputs to be used as training data. At the time of the experiments, these models ranked among the top-performing open-weight LLMs on the Chatbot Arena Leaderboard [^4] https://lmarena.ai/leaderboard (chiang2024chatbot).
The annotator models include the following: Llama 3.1 Nemotron 70B (wang2024helpsteer2), Qwen 2.5 72B (yang2024qwen2), Gemma 2 27B (gemma2024gemma), Command R+ 104B (cohere_for_ai_2024), and Mistral Large 2 123B [^5] https://mistral.ai/news/mistral-large-2407/. We apply Llama 3.3 70B (dubey2024llama) as the consolidator model. To address computational constraints, we use quantized versions available through the Ollama platform. [^6] https://ollama.com/ For synthetic data generation, we set the temperature to zero to obtain consistent results. For details on the models, see Appendix C.
6.2 Distillation
To produce OpeNLGaugeft, we apply LoRA hu2022lora with rank 16 and alpha 32 to fine-tune the instruction-tuned version of Llama 3.1 8B. The model is trained for one epoch with a learning rate of 2e-4, using AdamW optimizer (loshchilov2017weight) with a weight decay of 0.01. We apply a linear learning rate schedule with a warm-up period corresponding to the first 5% of training steps. The batch size is set to 16. This setup enables resource-efficient training that requires only around 20GB of VRAM, completing one training epoch in just six hours on a single A40 GPU.
6.3 Evaluation Datasets
We used seven popular meta-evaluation datasets to assess how our metric correlates with human judgments (for detailed descriptions of these datasets, including details on aggregation of human scores, see Appendix E). The datasets cover the following tasks: summarization – SummEval (fabbri2021summeval), QAGS (wang2020asking); story generation – HANNA (chhun2022human); data-to-text – SFRES and SFHOT (wen2015semantically), text simplification – Wiki-DA (alva2021suitability) and dialogue generation – TopicalChat (gopalakrishnan2023topical). For OpeNLGaugeft, text simplification is a task unseen during training and TopicalChat contains an unseen aspect (groundedness). For human evaluation of error spans, we used the RotoWire dataset (thomson2020gold) with annotations from a data-to-text task in the basketball domain, a task unseen during training by the evaluated fine-tuned metrics.
Dataset Llama 3.1 OpeNLGaugeft $\Delta$
QAGS 0.236 0.651 +0.415
SummEval 0.186 0.502 +0.316
TopicalChat 0.309 0.578 +0.269
SFRES/SFHOT 0.108 0.315 +0.207
HANNA 0.150 0.425 +0.275
Wiki-DA 0.405 0.789 +0.384
Table 4: Comparison of Spearman ($\rho$) correlations of the backbone model (Llama 3.1 8B) and our metric OpeNLGaugeft fine-tuned from the backbone on the dataset described in Section 5.1. For each dataset, the correlations are averaged across all evaluated aspects.
6.4 Baselines
We compare our methods with a variety of commonly used evaluation metrics, including traditional metrics such as ROUGE (lin-2004-rouge), BLEU (papineni2002bleu) and METEOR (agarwal2008meteor), distance-based metrics MoverScore (zhao-etal-2019-moverscore) and BERTScore (zhang2019bertscore), and trained metrics BARTScore (yuan2021bartscore) and UniEval (zhong2022towards). We also include the following LLM-based metrics: GPTScore (fu2024gptscore), G-Eval (liu-etal-2023-g), LLM Evaluation (chiang-lee-2023-closer), Prometheus (kim2023prometheus), Auto-J (li2023generative), InstructScore (xu-etal-2023-instructscore), TIGERScore (jiang2023tigerscore) and Themis (hu2024themis). To measure the improvement of OpeNLGaugeft relative to the base model, we include the instruction-tuned version of Llama 3.1 8B as an additional baseline. The specific metrics reported differ depending on the dataset and the evaluated NLG task.
Additionally, our comparisons include some task-specific and aspect-specific metrics: QAGS (wang2020asking) and FactCC (kryscinski-etal-2020-evaluating) for evaluating factual consistency in summarization, SARI (xu-etal-2016-optimizing) and LENS (maddela2022lens) for text simplification, and USR (mehri2020usr) for dialogue response generation tasks. The latter metric has several variants; we use the variant with the best Pearson correlation for each aspect in our comparison.
7 Results
Figure 3: Results of human evaluation of error spans and explanations. Top half of each bar: Error spans marked as correct or incorrect (hallucinated spans, no span provided, or spans without errors). Bottom half: Explanations marked as correct, partial (partially correct or incomplete) or incorrect (not addressing actual errors, vague or incorrect). The differences between TigerScore and OpeNLGaugeens are statistically significant (t-test, $p<0.05$). See Table 13 for more details.
Figure 4: Ablation results on QAGS and TopicalChat for OpeNLGaugeens. Plotted values represent differences in Spearman’s $\rho$ correlations with human scores between the ensemble with the original prompt and the corresponding ablation. For TopicalChat, Coh. = coherence, Eng. = engagingness, Gro. = groundedness, Nat. = naturalness, Avg. = average for all aspects.
7.1 Score Correlation with Humans
The results for factual consistency on QAGS are presented in Table 3 and the results for other datasets are available in Tables 1420 in Appendix 13. On QAGS, OpeNLGaugeens achieves the highest average performance on both Pearson’s $r$ and Spearman’s $\rho$. On Kendall’s $\tau$, our method is outperformed by Themis, which can be attributed to different score granularities used by these two methods (see Section G.1 for a discussion). The distilled version of our metric was consistently the third best measure. Notably, it outperformed the metrics based on the proprietary GPT-4 on this task.
On SummEval (Table 14), the best performing metric was Themis, closely followed by OpeNLGaugeens. However, note that training data for Themis include almost 62,000 human-annotated examples for the summarization task. Comparing our approach to TigerScore, another method that provides a similar level of explainability (error span annotations), we observe 15 p.p. improvements on Spearman’s $\rho$ averaged over all aspects. The smaller OpeNLGaugeft outperformed all other fine-tuned LLM-based metrics except Themis, and also surpassed metrics based on prompting GPT-3.5 by a large margin.
On TopicalChat (Table 15), the LLM Evaluation metric based on GPT-4 emerged as the strongest evaluator, while OpeNLGaugeens ranked between this method and its GPT-3.5-based version. OpeNLGaugeft achieved 27 p.p. improvement in Spearman’s $\rho$ compared to its Llama 3.1 backbone.
In data-to-text, OpeNLGaugeens excelled in evaluating naturalness, while OpeNLGaugeft stands out as the strongest evaluator of informativeness on the SFRES dataset (Table 16). Interestingly, on data-to-text problems, the distilled version of our metric achieved slightly better results on average than our ensemble. A similar situation is observed for story generation (Tables 1719), where the distilled version obtained better average scores on Spearman’s $\rho$ and Kendall’s $\tau$, and closely followed OpeNLGaugeens on Pearson’s $r$.
In text simplification (Table 20), our ensemble achieved superior performance across the board, outperforming LENS, a strong baseline metric specialized for this task.
Additionally, Table 4 presents a comparison of OpeNLGaugeft with Llama 3.1 8B (instruct), indicating a considerable improvement over the prompted model.
Generalization to unseen tasks
Although not originally trained on text simplification, OpeNLGaugeft outperforms all baseline metrics and individual LLMs in averaged Pearson correlation on the Wiki-DA dataset (Table 20) and achieves a particularly high score on meaning preservation.
Generalization to unseen aspects
On TopicalChat (Table 15), a large improvement for OpeNLGaugeft over Llama 3.1 8B is observed on groundedness, which is an aspect unseen during training. OpeNLGaugeft also surpassed most fine-tuned metrics and most individual LLMs on this aspect. Moreover, Wiki-DA contains additional unseen aspects (meaning preservation and simplicity) on which OpeNLGaugeft shows considerable improvement over Llama 3.1 8B and outperforms most of the individual larger LLMs.
7.2 Human Evaluation of Error Spans
We performed a small in-house human evaluation study to compare the quality of explanations obtained by OpeNLGauge and TigerScore, another LLM-based metric that also provides error-span annotations (see Table 1). For this purpose, we used a data-to-text task in the basketball domain (thomson2020gold). Five expert annotators evaluated the output of all three systems on 40 instances (a total of 120 outputs and 950 error spans). We asked the annotators to: (1) evaluate provided error spans, marking them as correctly identified, not containing an error, hallucinated, and situations where no span was provided; (2) evaluate generated explanations, marking them as correct, partially correct, incomplete, vague, incorrect, or texts that do not describe an error.
To assess reliability of our human annotation, we computed Cohen’s $\kappa$ coefficient (cohen1960coefficient) of inter-annotator agreement on 50 error spans with double annotations. We obtained $\kappa=0.82$ for the evaluation of error spans, and $\kappa=0.46$ for error explanations.
The results are shown in Figure 3, with more details provided in Table 13 in Appendix 13. Both OpeNLGaugeft and OpeNLGaugeens are over twice more accurate than TigerScore at annotating error spans, while finding over ten times more correct error spans. The task of providing accurate error explanations was more difficult for all the approaches evaluated. OpeNLGaugeens achieved the highest performance and was almost twice as accurate as TigerScore and OpeNLGaugeft, which achieved similar accuracies.
7.3 Ablation Experiments
Prompt ablations
We explore the effect of using different scales for the overall score and error severity: integer scales for both (Ablation 1), integer scale for overall score and categorical scale for severity (Ablation 2), and categorical scale for both (Ablation 3). Recall that OpeNLGauge uses a categorical scale for overall score and integer scale for error severity – see Section 4.
Correlation differences between the full prompt and the ablations are summarized in Figure 4, for detailed results see Appendix H.1. Overall, the change of scale has little effect on the average correlation of the whole ensemble, but it has a dramatic effect on some individual LLMs and aspects. This illustrates the ability of the ensemble to compensate for weaknesses in individual annotator models.
Finally, we examine the effect of removing the evaluation rules from the prompt (Ablation 4), which has an inconsistent effect on different models/aspects, but on average degrades the ensemble evaluation quality – up to 5.6 p.p.
Ensemble structure
We also analyze the effect of ensemble size and the influence of its particular components on the correlations with human scores by recomputing the results for all ensemble combinations. The results are presented in Appendix H.2. For Wiki-DA, performance increases with ensemble size, with the full ensemble being the best combination. For other datasets, there are a few smaller combinations that actually rank higher, but none is consistently better than the full ensemble.
Inter-annotator agreement between LLMs
To obtain additional insights into how the individual models of the ensemble diverge in their overall score predictions, we compute several measures of inter-annotator agreement. The results presented in Appendix I indicate only low to moderate agreement for most of the datasets, especially on exact overall scores, which suggests a sufficient diversity for combining outputs of these models into an ensemble.
Error analysis aggregation
The consolidator model aggregates error span annotations from multiple LLMs, which could potentially lead to an overall larger number of detected errors. To estimate the extent of over-annotation by both the ensemble and its components, we analyze the number of detected errors for output-aspect pairs rated with maximum score by human annotators. The results presented in Appendix K indeed show some tendency of the ensemble to over-annotate due to error accumulation from the individual models. However, most spans annotated by the ensemble were marked as correct in the experiment in Section 7.2. This could indicate that OpeNLGaugeens finds subtle errors which human annotators overlook, but further analysis of this discrepancy is needed.
Score aggregation
Additionally, we compare different methods of aggregating overall scores of individual LLMs to a final score in Appendix J. These results indicate that despite its simplicity, simple averaging is the most effective approach, generally providing the highest correlations with human scores.
8 Summary
In this work, we present OpeNLGauge – a versatile method for evaluating NLG that uses only open-weight models and provides fine-grained explainability. The method provides a much better explanation quality than previous methods and achieves competitive correlations with human judgments.
Limitations
OpeNLGauge is a method for evaluating a variety of NLG tasks. While this paper presents the evaluation on several NLG tasks and the method achieves good performance on unseen aspects, domains and tasks, the actual performance on new NLG tasks is unknown. In particular, the metric has not been tested in a multilingual setting. Moreover, previous research has shown that some LLM-based metrics have a bias towards texts generated by LLMs liu-etal-2023-g.
Acknowledgments
This work was supported by the European Research Council (Grant agreement No. 101039303, NG-NLG) and the National Recovery Plan funded project MPO 60273/24/21300/21000 CEDMO 2.0 NPO. It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
Appendix A Aspects and datasets in the training data
This section provides a detailed overview of all source datasets and evaluation aspects used to generate our training data for OpeNLGaugeft.
A.1 Summarization
As source data for summarization, we utilize five datasets spanning three distinct tasks: news article summarization, forum post summarization and dialogue summarization. Our evaluation dataset includes six commonly used aspects: four related to meaning of the evaluated text, one to its form and one to both. Some of these aspects overlap significantly in their definitions. This is intentional, as our goal is to make the evaluator model robust to small variations in aspect definitions, and enable generalization to new aspects. Table 6 provides an overview of all summarization systems in our dataset.
A.1.1 Datasets
  • CNN/DailyMail (hermann2015teaching) is a popular summarization dataset that consists of news articles from CNN and DailyMail, paired with corresponding bullet-point summaries. Originally developed for question answering, it was later adapted for summarization (nallapati2016abstractive). For this dataset, we use pre-generated outputs from stiennon2020learning, which include results from 11 different systems, including human references and an extractive baseline. Since our meta-evaluation datasets for summarization contain inputs from CNN/DailyMail, we ensure there is no overlap in the source texts between these datasets when sampling the inputs.
  • Newsroom (grusky2018newsroom) is a large-scale dataset of news articles and their summaries, collected from diverse sources, domains, authors and time range. In addition to newly generated LLM outputs, we use outputs of the systems evaluated in the original paper, which include three summarization systems and two extractive baselines.
  • SAMSum dataset (gliwa2019samsum) addresses the dialogue summarization task, containing dialogues with short summaries created by linguists. In addition to outputs from several different LLMs, we use pre-generated outputs of six systems from (gao2022dialsummeval).
  • TL;DR (volske2017tl) is a collection of Reddit posts with user-created summaries. Unlike many popular datasets that focus on news articles, TL;DR contains informal and less structured texts spanning diverse topics. Similarly to CNN/DailyMail, we use outputs from stiennon2020learning.
  • XSum (narayan2018don) is a dataset for the extreme summarization task, designed for abstractive approaches. Unlike datasets such as Newsroom or CNN/DailyMail, which favor extractive summarization, XSum contains BBC articles paired with concise, single-sentence summaries. We utilize pre-generated outputs from five systems and baselines evaluated in the original paper, along with newly generated output from LLMs.
A.1.2 Aspects
  • Consistency evaluates whether the summary is factually aligned with the source text. This involves determining if the facts in the summary can be entailed by the source. Consistency is closely tied to hallucinations, which may be categorized as factual or non-factual. In our approach, all information is required to be supported by the source text, making both types of hallucinations inconsistent with the source. The definition of consistency used in our prompts is: Extent to which the facts in the summary are consistent with the source text. Factually consistent summary should not contain facts that are not supported by the source text.
  • Accuracy is largely synonymous with consistency. It evaluates whether the factual information from the source is accurately represented in the summary. Following stiennon2020learning, we define accuracy as: Extent to which the factual information in the summary accurately matches the source text. An accurate summary should not contain information that is not present in the source text, should not contradict the source text, and generally should not be misleading.
  • Relevance of a summary is concerned with content selection. A relevant summary should include important points from the source text while omitting unimportant details. Compared to consistency, relevance is more subjective, as determining what information should or should not be included in a summary can sometimes be ambiguous. The definition used in our evaluation is: Extent to which the summary captures important information of the source text. A relevant summary should include all and only important information from the source text.
  • Coverage evaluates how much of the important information from the source text is covered by the summary. In this sense, it is closely related to relevance, however, it does not include non-redundancy as a criterion. We use a definition adapted from stiennon2020learning: Extent to which the summary covers the important information in the source text. A summary has good coverage if it mentions the main information from the source text that is important to understand the events described in the text. A summary has poor coverage if someone reading only the summary would be missing several important pieces of information about the event in the source text.
  • Coherence refers to the structural quality of the summary and involves attributes such as cohesion, consistency and relevance (reinhart1980conditions). It is determined by both semantic and formal structure of the text. We define coherence as: Extent to which the summary is well-structured and organized, presenting information in a logical order that flows naturally from sentence to sentence. Coherent summary forms a unified body of information and makes it easy to understand the main ideas.
  • Fluency focuses on the formal quality of the text, including grammaticality and naturalness. Unlike coherence, fluency is concerned with sentence–level quality rather than the overall structure of the text. We define fluency as: Formal quality of individual sentences of the summary. A fluent sentence should be grammatical, natural and easy to understand.
Format Example Datasets
JSON {"name": "The Phoenix", "eatType": "pub", "food": "Indian", ...} E2E NLG
[…]
attribute-value (2) name[The Phoenix], eatType[restaurant], food[Indian], ... E2E NLG
RDF (1) "The_Velvet_Underground | genre | Proto-punk" ... WebNLG 2020
RDF (2) (The_Velvet_Underground, genre, Proto-punk) ... WebNLG 2020
[…]
linearized table <page_title> List of Norwegian fjords </page_title> <section_title> ... ToTTo
Table 5: Input data formats used in our dataset for data-to-text tasks.
A.2 Data-to-text
For data-to-text category, we collected inputs from four datasets that represent four distinct tasks: table-to-text, RDF-to-text, attribute-value list to text and logical NLG. Since it is important for an NLG evaluation method to reliably evaluate outputs with respect to structured data in different formats, we include a number of different input formats in our training data, including JSON, CSV and linearized tables with markup. Table 5 provides an overview of the input formats. The list of evaluated systems in our dataset is shown in Table 7.
A.2.1 Datasets
  • E2E NLG dataset (duvsek2020evaluating; novikova2017e2e) was chosen as a representative of simple data-to-text tasks. In this dataset, models are tasked with generating descriptions of restaurant venues based on attribute-value-based meaning representations (MRs). Target descriptions were crowdsourced using textual and pictorial representations of MRs as stimuli. To ensure diversity, we selected pre-generated outputs from five systems described by (duvsek2020evaluating), considering their model architectures and evaluation results across various metrics and aspects. These were extended by new outputs from several LLMs. The inputs and pre-generated outputs were sampled from the test set. Three different input formats are used in the training data (see Table 5).
  • WebNLG 2020 (ferreira20202020) is an RDF-to-text dataset designed for generating natural language text from RDF triples collected from DBPedia knowledge base (mendes2012dbpedia). Each input contains between one and seven triples, where each triple represents a binary relation in the form (subject, property, object). Similarly to E2E NLG, we selected a diverse set of output systems based on model architectures and evaluation results from the WebNLG+ 2020 Challenge, and used additional LLMs to generate new outputs.
  • ToTTo (parikh2020totto) is an open-domain table-to-text generation dataset. It consists of Wikipedia tables with highlighted cells, and the task is to generate single-sentence descriptions of the data in these highlighted cells. Inputs are provided in two formats: full tables, which include indices to highlighted cells, and linearized tables, where only the highlighted data is presented in a linear order, while the structure is annotated with markup tags. As we observed that even medium-sized open-weight LLMs often struggle with hallucinations when using full tables, we restricted our evaluation to linearized tables.
  • LogicNLG (chen2020logical) introduced the task of logical NLG, where models generate statements that can be logically entailed from the data in a table. This task involves various aggregations and comparisons, making it more difficult than simply transforming structured data to free-form text. Although the task explicitly requires generating five logical statements per input, we sampled between one to five statements from each generated output to increase the diversity of output lengths. Inputs were formatted as CSV with “|” as a separator. In addition to new LLM outputs, we used existing system outputs from zhao-etal-2023-investigating.
A.2.2 Aspects
  • Faithfulness measures whether all information in the generated text is supported by the data, making it equivalent to precision. Similarly to factual consistency in summarization, we consider both factual and non-factual hallucinations as errors. For our evaluations, we define faithfulness as: Extent to which the information in the text is supported by the data.
  • Correctness evaluates whether the information from the data is accurately presented in the generated text. The output is maximally correct if it does not contain any incorrect statements with respect to the input. Correctness overlaps significantly with faithfulness, but its definition varies based on the specific task. For example, we define correctness for LogicNLG as: Extent to which the statements are logically and factually correct with respect to the provided data.
  • Coverage refers to the degree to which the generated text covers the information in the data. The output has maximum coverage when all information from the data is included in the text, which makes it analogous to recall. For example, in WebNLG 2020, it evaluates whether all predicates and their arguments are mentioned, while in ToTTo, it determines whether all highlighted table cells are described. We apply coverage to evaluate all datasets except LogicNLG, where the task is to infer interesting observations, rather than fully cover the source data. We define coverage as follows, with slight variations depending on the task: Extent to which the text includes description of all information presented in the data.
  • Informativeness is closely related to coverage, as it evaluates how much of the information the generated text provides. However, it does not require complete coverage of the data, therefore it is applicable to tasks such as LogicNLG, where full coverage of a table is not necessary. The definition used in our evaluation depends on the particular dataset. For example, the definition used for LogicNLG is: Extent to which the statements provide interesting or useful information about the data.
  • Fluency refers to the formal quality of the generated text, and includes grammaticality, naturalness and readability. Some definitions also include coherence (ferreira20202020), although coherence is usually treated as a separate aspect. For our evaluation, we define fluency as: Extent to which the text is grammatical, natural and easy to understand.
  • Grammaticality focuses on the correctness of grammar and spelling in the generated text. A text is fully grammatical if it contains no grammatical or spelling errors. While grammaticality is often included as a sub-aspect of fluency, both aspects are commonly used in practice. Therefore, we include it to help the model learn differences between evaluation aspects on different levels of hierarchy. In our dataset, grammaticality is defined as: Extent to which the text is grammatical (free of grammar and spelling errors).
  • Naturalness refers either to the human-likeness of the text, or the likelihood that it was produced by a native speaker. Like grammaticality, naturalness is often treated as a component of fluency. Additionally, its evaluation often includes assessment of grammaticality, as this can often be an indicator of whether the text was produced by a native speaker. This illustrates how evaluation aspects often overlap or have hierarchical relationships. For our purposes, naturalness is defined as: Extent to which the text is likely to have been produced by a native speaker.
A.3 Dialogue Response Generation
For dialogue response generation, we source the inputs from three dialogue datasets, focusing on open-domain non-task-oriented dialogue response generation. Table 8 provides an overview of all evaluated systems in the training dataset.
A.3.1 Datasets
  • Wizard of Wikipedia (dinan2018wizard) consists of conversations grounded in one of 1365 topics and corresponding knowledge retrieved from Wikipedia. In these conversations, either participant may select the topic and initiate the discussion, although they have asymmetric roles. One participant takes on the role of the wizard, an expert with access to a topic-relevant knowledge, on which they can base their responses. The other participant acts as an apprentice, a curious learner that is eager to discuss the chosen topic. To create inputs of varying length, we randomly select a dialogue history length between two turns and the full conversation, and truncate the dialogue to this length. The last utterance is replaced by a system output, except when the reference is used as evaluated output. In addition to newly generated LLM responses, we include human responses from the original dataset in the outputs.
  • EmpatheticDialogues (rashkin2018towards) is a dataset of dialogues grounded in emotional situations, designed to train and evaluate dialogue models on empathetic response generation. Each conversation is associated with an emotional label, where one of the participants describes a situation in which they experienced a given emotion. We sample up to five turns from each dialogue and replace the last utterance with a generated response. The emotion label is used as additional context for annotator LLMs to evaluate the appropriateness and empathy of the responses, but is excluded from the prompts used for system output generation.
  • DailyDialog (li2017dailydialog) includes conversations on various daily life topics, annotated with emotion labels and communicative intents. We include data from DailyDialog to represent diverse topics and scenarios in the training set. Alongside newly generated responses, we also collect pre-generated outputs from three sources (gupta2019investigating; huang2020grade; zhao2020designing) to represent older dialogue systems.
A.3.2 Aspects
  • Coherence in dialogue is a concept slightly different from coherence in tasks that involve generation of standalone texts, such as summarization or story generation. In dialogue, it measures how meaningful and logically consistent the response is with the preceding conversation. This includes not only alignment of the response with the last utterance, but also consistency with the dialogue participant’s earlier responses in terms of logic and style. In our evaluation, coherence is broadly defined as: Extent to which the response is a meaningful continuation of previous dialogue.
  • Relevance evaluates how closely a response aligns with the topic of conversation. In this sense, this aspect overlaps with coherence to some degree. We define relevance as: Extent to which the response is relevant and on-topic given the dialogue history.
  • Appropriateness addresses whether the response is semantically and pragmatically appropriate in the given context. Depending on the definitions, appropriateness might overlap to a large extent with coherence and relevance. For our evaluation, we define appropriateness as: Extent to which the response is semantically and pragmatically appropriate given the conversation history.
  • Empathy is evaluated specifically on responses from the EmpatheticDialogues dataset, where the goal is to determine whether the response acknowledges and reflects the emotions of the other participant. We define empathy as: Extent to which the response shows understanding of the feelings of the person talking about their experience.
  • Interestingness is concerned with the informational value of the response, specifically whether it presents stimulating ideas, facts or opinions. As it is one of the more subjective evaluation aspects, we define it vaguely and let the annotator models determine the criteria for interestingness: Extent to which the response is interesting given the dialogue history.
  • Engagingness is closely related to interestingness but is sometimes treated as a distinct aspect (e.g., mehri2020unsupervised; see-etal-2019-makes). While interestingness focuses on the context itself, engagingness emphasizes maintaining the user’s attention and encouraging them to continue with the conversation. For our purposes, engagingness is defined as: Extent to which the response captures and maintains the user’s interest, encouraging further interaction. Engaging responses contain opinions, preferences, thoughts or interesting facts.
  • Fluency in dialogue response generation has a similar meaning to its use in other tasks and refers to the formal quality of the response. We define fluency as: Extent to which the response is grammatically correct, natural and fluent.
  • Understandability evaluates both the content and form of a response, focusing on its clarity and ease of comprehension. The definition we use for our evaluation is: Extent to which the response is easy to understand and comprehend given the dialogue history.
A.4 Story Generation
The NLG tasks discussed so far generally contain inputs that are relatively longer compared to the outputs. This pattern is especially common in tasks like summarization and dialogue generation, although certain data-to-text tasks also share this characteristic. To represent scenarios with short inputs and long outputs, we include story generation in the training data. Table 9 lists all evaluated systems in our dataset.
A.4.1 Datasets
As the source of inputs, we use WritingPrompts (fan2018hierarchical), a story generation dataset derived from Reddit’s WritingPrompts subreddit, where users submit prompts that can inspire other users to write stories. The dataset consists of a diverse range of topics, story lengths and writing styles. We reuse existing outputs from the OpenMEVA dataset (guan2021openmeva) and generate additional outputs by four LLMs. To increase the diversity of generated stories in terms of their length, we generate the outputs with two different prompt versions, each requiring a different length of the story. The outputs are then randomly sampled from either the shorter or the longer set. As we observed a tendency of LLMs to generate a long list of errors for longer inputs, our evaluator models are instructed to limit the number of identified errors to a maximum of eight, and to prioritize the most severe ones if necessary.
A.4.2 Aspects
While relevance and coherence are two commonly used aspects for story generation, there is no consensus on which other evaluation aspects are the most relevant. Inspired by social sciences, chhun2022human propose four additional aspects, aimed at providing a complete and non-redundant set of criteria. Following their work, we adopt the aspects defined in the HANNA benchmark, using our own definitions for most of them:
  • Relevance measures the degree to which a story aligns with the given prompt (chhun2022human; chiang2023can), title (jhamtani2020narrative; yao2019plan; xie2023next) or story beginning wang2020narrative. In some cases, relevance also evaluates whether the story remains on-topic for its duration (e.g., goldfarb2020content). Since our inputs are prompts, we define relevance as: Extent to which the story is relevant to the writing prompt.
  • Coherence in story generation typically refers to logical consistency and narrative flow (e.g., yao2019plan; li2023collaborative; jhamtani2020narrative). Other works define coherence more vaguely, such as how much the story “makes sense” (chhun2022human) or how well its sentences “fit together” (xie2023next). For our purposes, coherence is defined as: Extent to which the story is logically consistent and coherent.
  • Engagement is a subjective and often vaguely defined aspect that evaluates how engaging the story is to the reader (e.g., chhun2022human; li2023collaborative). Due to its inherent subjectivity, we apply a simple definition and leave its interpretation to the evaluator models: Extent to which the story is engaging and interesting.
  • Empathy is related to emotional commentary and empathy, and refers to how well the story conveys character’s emotions. We define empathy as: The clarity and depth with which the character’s emotions are conveyed in the story.
  • Surprise is concerned with the story’s ending, and evaluates its unexpectedness and originality. We define surprise as: How surprising the end of the story was.
  • Complexity measures how intricate and elaborate the story is. Complexity is not necessarily an aspect of quality, but rather a feature of the text. Whether greater complexity is desired or not depends on the audience. Our definition of complexity is: How elaborate the story is.
A.5 Question Answering
The question answering subset of our dataset consists of two distinct tasks: narrative question answering and table question answering. The inputs consist of a question and structured data in which the answer should be grounded. The evaluated systems in our training data are listed in Table 10.
A.5.1 Datasets
  • NarrativeQA (kovcisky2018narrativeqa) consists of human-written questions and free-form answers based on stories or their summaries. The stories include books and movie scripts and are provided either as full texts or as human-written summaries. Since full stories do not fit into the context window of many LLMs, we use only summaries to generate the outputs. Along with newly generated answers, we include human reference answers from the original datasets in the evaluated outputs.
  • FeTaQA (nan2022fetaqa) is a question answering dataset based on Wikipedia tables that requires models to aggregate and reason about the entities in the table and their relations. This places the task at the intersection of data-to-text and question answering task categories. In addition to new LLM outputs, we use pre-generated outputs from zhao-etal-2023-investigating.
A.5.2 Aspects
  • Correctness evaluates if the answer to a question is correct with respect to the input. Since our models are instructed to assess the quality on an ordinal scale, we evaluate a degree of correctness – the answer should receive the maximum score if it is fully correct, while lower scores should reflect the number and severity of correctness issues. We define correctness as: Extent to which the answer to the question is correct with respect to the input.
  • Informativeness addresses whether all information required by the question is provided in the answer. We define informativeness as: Extent to which the answer provides all information that the question asked for.
  • Completeness evaluates comprehensiveness of the answer and the degree to which all aspects of the question are covered. The meaning is slightly different from informativeness, which is concerned with the information that the question explicitly asks for. In our dataset, completeness is defined as: Extent to which the answer is comprehensive and ensures all question aspects are addressed.
  • Conciseness measures the degree to which an answer is focused and directly answers the question without unnecessary details and elaboration. Although the goal might often be to generate both complete and concise answers, these two aspects might correlate negatively. Conciseness is defined as: Extent to which the answer is concise and to the point.
  • Relevance is concerned with the specificity of an answer and measures the degree to which the answer addresses the particular question asked. Although it is related to conciseness, relevance is not that much concerned with the amount of detail in the answer. We define relevance as: Extent to which the answer is specific and meaningful with respect to the question.
  • Factuality evaluates factual consistency of the answer with the provided context. In our dataset, this aspect is used in the narrative question answering task, and we use a similar definition as in summarization: Extent to which the answer is supported by the summary.
  • Faithfulness is used for the table question answering task and is synonymous with factuality. We define faithfulness as: Extent to which the information presented in the answer is supported by the input.
  • Fluency in question answering refers to the formal quality of the answer and is defined similarly as in the other tasks presented so far: Extent to which the response is grammatical, natural and easy to understand.
  • Naturalness is interpreted in the same way as in data-to-text and dialogue response generation tasks, and the definition we apply is: Extent to which the answer is likely to have been produced by a native speaker.
  • Grammaticality measures the grammatical quality of the answer and is defined as: Extent to which the answer is grammatical (free of grammar and spelling errors).
Appendix B Collected system outputs
Tables 610 list the evaluated systems for each task category in our dataset.
Table 6: Overview of the evaluated systems in the dataset for the summarization task. Sources: GW = gao2022dialsummeval, GR = grusky2018newsroom, NA = narayan2018don, ST = stiennon2020learning, new = newly generated.
System Type Sources
Qwen 2.5 0.5B (yang2024qwen2) Instruction-tuned LLM new
Llama 2 7B Chat (touvron2023llama) Instruction-tuned LLM new
Gemma 2 2B (gemma2024gemma) Instruction-tuned LLM new
Nous Hermes 2 Mixtral 8x7B DPO [^7] https://huggingface.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-DPO Instruction-tuned LLM new
GPT-4o [^8] https://openai.com/index/hello-gpt-4o/ Instruction-tuned LLM new
OpenAI summarization (pre-trained) pre-trained LMs ST
OpenAI summarization (supervised) LMs trained with SFT ST
OpenAI summarization (RLHF) LMs trained with SFT+PPO ST
T5 (raffel2020exploring) pre-trained LM ST
UniLM (dong2019unified) pre-trained LM ST
CODS (wu-etal-2021-controllable) BART-based hybrid model GW
ConvoSumm (fabbri-etal-2021-convosumm) BART-based model GW
Ctrl-DiaSumm (liu-chen-2021-controllable) BART-based model GW
ConvS2S (gehring2017convolutional) Convolutional seq-to-seq NA
Topic-ConvS2S (narayan2018don) Topic-conditioned ConvS2S NA
S2S (cho2014learning; sutskever2014sequence) RNN-based seq-to-seq with attention GR
PNG (vinyals2015pointer; gu̇lccehre2016pointing) Pointer-generator network GR, GW, NA
TextRank (barrios2016variations) Ranking-based extractive summarization GR
Reference Human-written reference ST
Title Extractive baseline ST
LEAD Extractive baseline NA
LEAD-2 Extractive baseline ST
LEAD-3 (see2017get) Extractive baseline GR, GW, ST
Ext-Oracle Extractive oracle NA
Fragments Extractive oracle NA
Table 7: Overview of the evaluated systems in the dataset for the data-to-text task. Sources: DU = duvsek2020evaluating, FE = ferreira20202020, ZH = zhao-etal-2023-investigating, new = newly generated.
System Type Sources
Qwen 2.5 Coder 1.5B (yang2024qwen2) Instruction-tuned LLM new
Gemma 2 2B (gemma2024gemma) Instruction-tuned LLM new
Llama 2 7B Chat (touvron2023llama) Instruction-tuned LLM new
Solar 10.7B (kim-etal-2024-solar) Instruction-tuned LLM new
DeepSeek Coder v2 16B (zhu2024deepseek) Instruction-tuned LLM new
Nous Hermes 2 Mixtral 8x7B DPO [^9] https://huggingface.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-DPO Instruction-tuned LLM new
Claude 3.5 Sonnet [^10] https://www.anthropic.com/news/claude-3-5-sonnet Instruction-tuned LLM new
GPT-4o [^11] https://openai.com/index/hello-gpt-4o/ Instruction-tuned LLM new
NILC (sobrevilla-cabezudo-pardo-2020-nilc-webnlg) Fine-tuned BART FE
Orange-NLG (montella2020denoising) Fine-tuned BART FE
Amazaon AI (Shanghai) (guo-etal-2020-2) Graph CNN + T5 FE
GPT2-C2F (chen2020logical) Fine-tuned GPT-2 ZH
LoFT (zhao2023loft) Fine-tuned BART ZH
PLOG (liu-etal-2022-plog) Fine-tuned T5 ZH
R2D2 (nan2022r2d2) Fine-tuned T5 ZH
Flan-T5 (chung2024scaling) Instruction-tuned LM ZH
Adapt (elder-etal-2018-e2e) RNN seq-to-seq DU
Sheff2 (chen2018sheffield) RNN seq-to-seq DU
Slug (juraska-etal-2018-deep) RNN + convolutional seq-to-seq DU
Forge1 (mille2018forge) Rule-based DU
TR2 (smiley-etal-2018-e2e) Template-based DU
DANGNT-SGU (tran-nguyen-2020-webnlg) Template-based FE
RALI-Université de Montréal (lapalme-2020-rdfjsrealb) Template-based FE
Table 8: Overview of the evaluated systems in the dataset for the dialogue response generation task. Sources: GU = gupta2019investigating, HU = huang2020grade, ZH = zhao2020designing, new = newly generated.
System Type Sources
Claude 3.5 Sonnet [^12] https://www.anthropic.com/news/claude-3-5-sonnet Instruction-tuned LLM new
GPT-4o [^13] https://openai.com/index/hello-gpt-4o/ Instruction-tuned LLM new
Tülu 3 (lambert2024tulu) Instruction-tuned LLM new
Dolphin 2.9 Llama 3 8B [^14] https://huggingface.co/cognitivecomputations/dolphin-2.9-llama3-8b Instruction-tuned LLM new
Vicuna 7B (zheng2023judging) Instruction-tuned LLM new
BlenderBot-small (roller-etal-2021-recipes) Dialogue LM new
DialoGPT-small (zhang-etal-2020-dialogpt) Dialogue LM new
GPT-Neo 125M (gao2020pile) Pre-trained LM new
GPT-2 (wolf2019transfertransfo) Pre-trained LM ZH
CVAE (zhao2017learning) Conditional variational autoencoder GU
HRED (serban2016building) Hierarchical recurrent encoder-decoder GU, ZH
Transformer-generator (dinan2018wizard) Transformer-based generative model HU
Transformer-ranker (urbanek-etal-2019-learning) Transformer-based ranking model HU
DualEncoder (lowe2015ubuntu) LSTM dual encoder GU
VHRED (serban2017hierarchical) Latent variable HRED ZH
S2S (cho2014learning; sutskever2014sequence) RNN-based seq-to-seq with attention GU, ZH
Table 9: Overview of the evaluated systems in the dataset for the story generation task. Sources: GU = guan2021openmeva, new = newly generated.
System Type Sources
Gemma 2 2B (gemma2024gemma) Instruction-tuned LLM new
Dolphin 2.9 Llama 3 8B [^15] https://huggingface.co/cognitivecomputations/dolphin-2.9-llama3-8b Instruction-tuned LLM new
Nous Hermes 2 Mixtral 8x7B DPO [^16] https://huggingface.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-DPO Instruction-tuned LLM new
GPT-4o [^17] https://openai.com/index/hello-gpt-4o/ Instruction-tuned LLM new
GPT-2 (radford2019language) Pre-trained LM GU
GPT-KG (guan2020knowledge) Knowledge-enhanced GPT-2 GU
Fusion (fan2018hierarchical) Convolutional seq-to-seq with attention GU
Plan&Write (yao2019plan) Hierarchical RNN-based model GU
S2S (cho2014learning; sutskever2014sequence) RNN-based seq-to-seq GU
Table 10: Overview of the evaluated systems in the dataset for the question answering task. Sources: KO = kovcisky2018narrativeqa, ZH = zhao-etal-2023-investigating, new = newly generated.
System Type Sources
Claude 3.5 Sonnet [^18] https://www.anthropic.com/news/claude-3-5-sonnet Instruction-tuned LLM new
GPT-4o [^19] https://openai.com/index/hello-gpt-4o/ Instruction-tuned LLM new
Nous Hermes 2 Mixtral 8x7B DPO [^20] https://huggingface.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-DPO Instruction-tuned LLM new
DeepSeek Coder v2 16B (zhu2024deepseek) Instruction-tuned LLM new
Llama 2 7B Chat (touvron2023llama) Instruction-tuned LLM new
Qwen 2.5 Coder 1.5B (yang2024qwen2) Instruction-tuned LLM new
Qwen 2.5 0.5B (yang2024qwen2) Instruction-tuned LLM new
GPT-Neo 125M (gao2020pile) Pre-trained LM new
BART (lewis2019bart) Pre-trained LM ZH
Flan-T5 (chung2024scaling) Instruction-tuned LM ZH
OmniTab (jiang-etal-2022-omnitab) Table pre-trained LM ZH
ReasTAP (zhao-etal-2022-reastap) Table pre-trained LM ZH
TAPEX (liu2022tapex) Table pre-trained LM ZH
Reference Human-written reference KO
Appendix C LLMs used in the experiments
The specific versions of LLMs used in the experiments are presented in Table 11.
Model Quantization Tag
Command R+ 104B 5-bit command-r-plus:104b-08-2024-q5_K_M
Gemma 2 27B 8-bit gemma2:27b-instruct-q8_0
Llama 3.1 Nemotron 70B 8-bit nemotron:70b-instruct-q8_0
Llama 3.3 70B 8-bit llama3.3:70b-instruct-q8_0
Mistral Large 2 123B 4-bit mistral-large:123b-instruct-2407-q4_K_M
Qwen 2.5 72B 8-bit qwen2.5:72b-instruct-q8_0
Table 11: Quantization levels and Ollama tags used for the models.
Appendix D Outliers
Table 12 shows the percentages of outliers detected for each annotator model in the ensemble and task category during synthetic data generation. To keep the merged evaluation outputs internally consistent, these outliers were removed before merging the outputs of the individual LLMs.
The proportions in the table indicate that the final annotation utilized most of the generated evaluations from the individual LLMs. The exceptions include Command R+ 104B and Gemma 2 27B on data-to-text tasks, with 29% and 10% outliers, respectively. Additionally, 10% evaluation outputs of Command R+ 104B were removed for question answering tasks. However, on average only 3.4% of evaluation outputs are detected as outliers across all model-task pairs.
Model Data-to-text Summarization Dialogue Story Generation Question Answering
Command R+ 104B 28.88 0.13 0.73 0.94 9.61
Gemma 2 27B 10.32 1.96 1.52 1.03 5.07
Llama 3.1 Nemotron 70B 4.27 1.48 4.02 1.03 0.85
Mistral Large 2 123B 4.89 0.93 0.63 0.25 1.22
Qwen 2.5 72B 1.39 1.36 1.43 1.12 0.47
Table 12: Proportions (%) of evaluation outputs detected as outliers for each model and task category. Across all model-task pairs, 3.4% of evaluation outputs are detected as outliers on average (median = 1.36%).
Appendix E Meta-evaluation datasets
We used the following datasets for meta-evaluation. In the datasets where human scores are not already aggregated, we averaged the scores of all annotators.
  • SummEval (fabbri2021summeval) is a standard meta-evaluation dataset for summarization. It consists of summaries generated from CNN/DailyMail articles, with 100 input articles and 16 different system outputs for each article. Human evaluations address four aspects: (factual) consistency, relevance, coherence, and fluency. Each output is scored by three expert annotators and five crowdworkers. We use only expert scores to measure correlations.
  • QAGS (wang2020asking) contains annotations of consistency for summaries from CNN/DailyMail and XSum datasets. It includes 235 CNN/DailyMail summaries and 239 XSum summaries, each annotated by three evaluators. Annotators assign a binary factual consistency score (yes/no) for each sentence of the summary. We follow (wang2020asking) and apply majority voting for each sentence annotation, followed by averaging sentence-level scores to obtain the overall score for the summary.
  • HANNA (chhun2022human) is a benchmark for story generation based on the WritingPrompts dataset. It contains three human scores for each of the 1056 stories across six aspects: relevance, coherence, engagement, empathy, surprise and complexity (see Appendix A.4 for details on these aspects).
  • SFRES and SFHOT (wen2015semantically) are used to perform a meta-evaluation for data-to-text. These datasets consist of dialogue acts (DAs) in structured format with generated responses providing information about restaurants and hotels in San Francisco. Human judgments are provided for informativeness and naturalness.
  • TopicalChat (gopalakrishnan2023topical) annotations from the USR dataset (mehri2020usr) are used for meta-evaluation of dialogue response generation. The dataset includes human evaluations for five aspects: groundedness, coherence, interestingness, naturalness and understandability. As the source data differ from our training data, TopicalChat serves as an out-of-domain evaluation dataset. Additionally, since groundedness is not present in our training data, we evaluate it as an unseen aspect.
  • Wiki-DA (alva2021suitability) is a dataset of DA human ratings for text simplification, a task unseen by OpeNLGaugeft during training. Along with scores for fluency, this dataset also includes two unseen aspects: meaning preservation and simplicity.
Appendix F Prompt templates
Prompt templates for annotator and consolidator LLMs are presented in Listings 12. Prompt template used for fine-tuning and inference of the distilled model is shown in Listing 3.
Appendix G Full results
System
Error span
Error
No span given
Not an error
Hallucination
Total
Explanation
OpeNLGaugeens
Correct 218 9 3 0 230
Partially correct 41 0 0 1 42
Incomplete 20 0 1 0 21
Vague 4 2 1 0 7
Incorrect 20 3 32 1 56
Not an error 0 0 10 0 10
Total 303 14 47 2 366
Span OK (%) 83
Exp. correct (%) 63
OpeNLGaugeft
Correct 108 3 4 4 119
Partially correct 60 1 2 0 63
Incomplete 9 0 0 0 9
Vague 3 9 0 0 12
Incorrect 41 7 42 6 96
Not an error 0 1 8 0 9
Total 221 21 56 10 308
Span OK (%) 72
Exp. correct (%) 39
TigerScore
Correct 8 17 0 0 25
Partially correct 5 2 0 0 7
Incomplete 2 2 0 0 4
Vague 0 3 0 0 3
Incorrect 5 16 2 0 23
Not an error 1 2 0 0 3
Total 21 42 2 0 65
Span OK (%) 32
Exp. correct (%) 38
Table 13: Detailed results of human evaluation of error-spans and their explanations. Each table shows absolute occurrence counts of different error span and explanation validity (see Section 7.2), with overall proportions of correct annotation given below. The reported differences between TigerScore and OpeNLGaugeens are statistically significant (t-test, $p<0.05$).
Tables 1420 contain the meta-evaluation results for additional datasets described in Appendix E. Detailed results of human evaluation of error span quality are presented in Table 13.
G.1 Kendall’s $\tau$ correlations
On QAGS, our method achieves lower Kendall’s $\tau$ than Themis, although it surpasses the metric on Spearman’s $\rho$. This discrepancy can be attributed to a difference in score precisions between these two methods. As a result of averaging, OpeNLGaugeens provides more granular floating point scores, while Themis predicts integer scores on a Likert scale. Generally, we can expect more tied pairs (i.e. pairs that are neither concordant nor discordant) in the calculation of Kendall’s $\tau$ with integer scores, which can have a substantial effect on the correlations. In QAGS, 78% of human scores map to integer scores after normalization. Therefore, we consider Spearman’s $\rho$ more appropriate for comparing evaluation capabilities of different metrics.
Similarly, specific individual LLMs score higher on Kendall’s $\tau$ than the ensemble on some datasets (see Tables 3, 14, 18 and 19). This is due to the same reason outlined above. Specifically, there are 37% tied pairs for OpeNLGaugeens and human scores, while the individual LLMs, which provide integer scores, show substantially larger proportions of tied pairs: between 46% and 54%.
Appendix H Ablation experiments
H.1 Prompt ablations
The detailed results of prompt ablation experiments are presented in Figure 1114.
H.2 Ensemble size ablations
The effect of the ensemble size on Spearman correlations are presented in Figures 510.
Appendix I Inter-annotator agreement between ensemble models
To obtain additional insights into the variance between individual ensemble LLMs in their predicted overall scores, we compute pairwise inter-annotator agreements between the models on all meta-evaluation datasets.
Figure 16 shows the Spearman correlations for each LLM pair and dataset used in our meta-evaluation, where the correlations are calculated over all evaluation aspects in a given dataset. The agreement is generally high on the QAGS and Wiki-DA datasets, although substantially lower on other datasets. Note that individual models also achieved relatively high correlations with humans on these two datasets, which indicates that they might generally consist of examples (and possibly evaluation aspects) for which it is easier for LLMs to make decisions on their quality. In contrast, the models show relatively low agreement on HANNA, which could be explained by the more subjective nature of the story generation task and the corresponding evaluation aspects. [^21] While such hypotheses could in principle be tested by comparing our results with the agreement between human annotators on a particular task, there are several issues that limit the reliability of such comparisons – for each meta-evaluation dataset we use, at least two of the following hold: (1) the number of human annotators is different from the number of LLMs in our ensemble, (2) either the sets of human annotators differ between evaluated outputs, or it is not made explicit in the dataset (or the corresponding paper) whether this is the case, (3) rating scales are different from ours (often three levels, or even binary). Across all tasks, Command R+ 104B tends to disagree more with other LLMs when compared to other pairwise agreements.
To assess the agreement of the models on exact overall score predictions, we also measure pairwise Cohen’s $\kappa$ (Figure 17), which treats the overall scores as categorical and therefore serves as a stricter measure. As expected, the agreements are lower compared to those measured by Spearman correlations, while the general trend is the same: the agreement on exact scores is higher on QAGS and Wiki-DA than on other datasets, while Command R+ 104B generally disagrees more with other LLMs across tasks.
Finally, we compute Krippendorff’s $\alpha$ (krippendorff2011computing), which allows us to measure the agreement between all ensemble models, while also being applicable to ordinal data. In general, the agreements are low to moderate. Similarly to the pairwise agreements, we observe substantially higher agreements on QAGS and Wiki-DA, while the models agree the least on scoring generated stories in the HANNA dataset.
Overall, our analysis indicates that the predictions of individual models are sufficiently diverse to benefit from the combination in an ensemble without much redundancy.
Appendix J Score aggregation methods
Table 21 compares results of different approaches to aggregation of ensemble scores:
  • Average computes the final score for the evaluated output as a simple average of scores from all individual LLMs.
  • Average w/o outliers first removes the outliers before computing the average. The score is considered an outlier if it differs from the average of other scores for the same example by at least two standard deviations and this difference is at least 1.
  • Median computes the final score as a median of the scores from individual LLMs to disregard the effect of extreme scores.
  • Majority voting selects the most frequent score from the individual LLMs for the given output.
  • Min selects the minimum of individual scores as the final score. This corresponds to the most strict evaluation, i.e. typically the evaluation that detected the most errors or assigned highest severity levels to the identified errors.
Appendix K False positives analysis
To estimate the extent of potential over-annotation by OpeNLGaugeens and its components, we analyze overall score predictions for output-aspect pairs $(y,a)$ which obtained maximum scores by all human annotators in a given dataset. We denote these examples $Y_{max}$ and assume that if $(y,a)\in Y_{max}$, then $y$ does not contain any errors with respect to aspect $a$.
[…]
Figure 18 shows the distribution of error counts (top), mean severity levels (middle) and maximum severity levels (bottom) per output, as predicted by OpeNLGaugeens and its components for the $Y_{max}$ subset of SummEval. Most individual LLMs assess the outputs in $Y_{max}$ as error-free, particularly Command R+ 104B, which agrees almost perfectly with human annotators in this subset. Note that in general, Command R+ 104B tends to disagree the most with other LLMs, as discussed in Appendix I, while also achieving the highest correlations with humans in evaluating factual consistency (Table 14), which represents approximately half of the examples in $Y_{max}$. In contrast, the ensemble shows a tendency to include at least one error in most of its annotations for $Y_{max}$. This could be attributed to the merging procedure by the consolidator model, which aggregates errors from five different models and could lead to accumulation of errors. As expected, an increasing number of errors detected by a model is also reflected in the mean and maximum severity levels.
The error annotations for TopicalChat (Figure 19) show a similar pattern, although a larger proportion of individual models have a tendency to detect one or more errors in this case.
Appendix L Output examples
Figures 2022 show additional output examples for RDF-to-text, dialogue response generation, and summarization tasks.
### Instructions
Your task is to evaluate an output of data-to-text task, where the model was instructed to write a single-paragraph description of a venue based on the given data. The data consist of a set of attribute-value pairs in the form attribute: value’.
Based on the given data and the generated text, identify errors in the text with respect to {{ aspect_name }} (described below).
For each error, determine its severity on a scale from 1 to 5, where 1 is the least severe and 5 is the most severe.
Definition of {{ aspect_name }}:
{{ aspect_definition }}
Rules:
Do not make assumptions and do not bring in external knowledge not present in the provided context.
Identify only the errors related to the {{ aspect_name }} of the text. Do not consider other aspects like {{ negative_aspects }}!
If there are no errors related to {{ aspect_name }} in the text, you should output No Error and provide Excellent score.
Steps:
1. Carefully read the data and identify the main attributes and their values.
2. Read the generated text and compare it with the source data with respect to {{ aspect_name }}.
3. If the text contains any error that negatively affects its {{ aspect_name }}, identify its exact location (specific word or phrase), explain why it is considered an error, and determine the severity of the error.
4. Finally, provide an overall score for the {{ aspect_name }} of the text. The score should be a label on the following scale (lowest to highest): Unacceptable’, Poor’, Fair’, Good’, Excellent’. The score Unacceptable indicates that the text is {{ min_score_desc }}, while Excellent indicates that the text is {{ max_score_desc }}.
### Data
{{ input }}
### Generated Text
{{ output }}
### Output format:
Generate your output exactly in this format:
Error 1:
Location: <location of the error - the exact word or phrase in the response>
Explanation: <explanation for the error, including the reason why it is considered {{ aspect_name }} issue>
Severity: <integer from 1 to 5>
Error 2:
Overall score: <one of: Unacceptable, Poor, Fair, Good, Excellent>
Explanation of the score: <explanation of the score>
Listing 1: Annotator prompt template for the data-to-text task.
### Instructions
You are given multiple error annotation sets for an AI model output. Your task is to merge the annotation sets to a single final annotation set.
The result shouldnt contain any duplicates. If there are multiple error annotations for approximately the same location that describe the same issue, you should merge them into single location. Otherwise, the error annotations should be as granular as possible. If there are multiple different locations with the same issue, each should have its own error annotation. Likewise, if there are multiple issues with respect to the same location, each should have its own error annotation. Use the following guidelines:
* Each error annotation should describe a single issue.
* Merge only annotations where the locations have significant overlap.
* When merging multiple locations, choose a single span from the output text that covers the locations from merged annotations.
* Never include multiple spans from the annotations under the same "Location" line.
* Do not include any other text in "Location" line than the text that is actually in the output, except for annotations that mention omissions or similar issues.
* When merging explanations, combine the most relevant information from the merged annotations.
* Severity levels range from 1 to 5 from least severe to most severe. Use the most severe level from the merged annotations.
* Final annotation set should not include more than 8 error annotations. If there are more than that, use only the most severe ones.
* Make the final annotation set as concise as possible in terms of number of error annotations.
Dont use any markdown formatting. Generate merged error annotations in this format, without any additional text:
Error 1:
Location: <span of text from the output, or None if not applicable>
Explanation: <explanation>
Severity: <severity level>
### Model output
{{ output }}
### Error annotations
{{ annotations }}
Listing 2: Consolidator prompt template for merging of individual annotations.
### Task
Your task is to evaluate a model output for {{ task_name }} task with respect to {{ aspect_name }}. {{ extra_task_info }}
### Aspect Definition
{{ aspect_name }} - {{ aspect_definition }}
### Dialogue history
{{ input }}
{% if context %}
### Knowledge
{{ context }}
{% endif %}
### Response
{{ output }}
### Instructions
For any error in the output, identify its location, assign a severity level and provide an explanation. Report at most 8 errors. If there are more errors, report only the most severe ones. Finally, provide an overall score between 0 and 100 for {{ aspect_name }} of the output.
Listing 3: Prompt template for the fine-tuned OpeNLGaugeft metric.
Metric Consistency Coherence Relevance Fluency Average
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
ROUGE-1 0.167 0.126 0.160 0.130 0.115 0.094 0.326 0.252 0.192 0.150
ROUGE-2 0.184 0.139 0.187 0.155 0.159 0.128 0.290 0.219 0.205 0.161
ROUGE-L 0.128 0.099 0.115 0.092 0.105 0.084 0.311 0.237 0.165 0.128
BERTScore 0.151 0.122 0.285 0.220 0.302 0.232 0.186 0.154 0.231 0.182
MOVERSscore 0.159 0.118 0.157 0.127 0.129 0.105 0.318 0.244 0.191 0.148
BARTScore 0.266 0.220 0.474 0.367 0.318 0.243 0.258 0.214 0.329 0.261
UniEval 0.446 0.371 0.575 0.442 0.426 0.325 0.449 0.371 0.474 0.377
G-Eval (GPT-3.5) 0.386 0.318 0.440 0.335 0.385 0.293 0.424 0.347 0.409 0.323
G-Eval (GPT-4) 0.507 0.425 0.582 0.457 0.548 0.433 0.455 0.378 0.523 0.423
LLM Evaluation (GPT-3.5) 0.393 0.331 0.459 0.371 0.455 0.363 0.355 0.296 0.415 0.340
LLM Evaluation (GPT-4) 0.531 0.464 0.540 0.434 0.491 0.395 0.480 0.409 0.511 0.426
X-Eval 0.428 0.340 0.530 0.382 0.500 0.361 0.461 0.365 0.480 0.362
Prometheus 0.150 0.137 0.150 0.126 0.164 0.138 0.189 0.168 0.163 0.142
AUTO-J 0.131 0.121 0.245 0.203 0.262 0.222 0.154 0.141 0.198 0.172
TIGERScore 0.427 0.387 0.381 0.318 0.366 0.304 0.363 0.327 0.384 0.334
InstructScore 0.232 0.213 0.328 0.276 0.211 0.179 0.260 0.237 0.258 0.226
Themis 0.600 0.566 0.566 0.485 0.474 0.412 0.571 0.533 0.553 0.499
OpeNLGaugeens 0.548 0.470 0.604 0.462 0.513 0.389 0.470 0.389 0.534 0.427
• Command R+ 104B 0.633 0.603 0.239 0.203 0.360 0.302 0.347 0.320 0.395 0.357
• Gemma 2 27B 0.459 0.421 0.455 0.386 0.428 0.358 0.427 0.386 0.442 0.388
• Llama 3.1 Nemotron 70B 0.559 0.517 0.469 0.391 0.419 0.356 0.358 0.325 0.451 0.397
• Mistral Large 2 123B 0.627 0.590 0.528 0.434 0.456 0.375 0.398 0.359 0.502 0.439
• Qwen 2.5 72B 0.567 0.521 0.525 0.433 0.388 0.317 0.433 0.389 0.478 0.415
Llama 3.1 8B 0.181 0.165 0.176 0.150 0.156 0.127 0.232 0.218 0.186 0.165
OpeNLGaugeft 0.527 0.453 0.561 0.441 0.514 0.408 0.407 0.349 0.502 0.413
Table 14: Segment-level Spearman ($\rho$) and Kendall ($\tau$) correlations of different metrics on SummEval.
Metric Groundedness Coherence Engagingness Naturalness Average
$r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$
ROUGE-L 0.193 0.203 0.176 0.146 0.295 0.300 0.310 0.327 0.243 0.244
BLEU-4 0.131 0.235 0.180 0.175 0.232 0.316 0.213 0.310 0.189 0.259
METEOR 0.250 0.302 0.212 0.191 0.367 0.439 0.333 0.391 0.290 0.331
BERTScore 0.214 0.233 0.226 0.209 0.317 0.335 0.291 0.317 0.262 0.273
USR 0.416 0.377 0.337 0.325 0.456 0.465 0.222 0.447 0.358 0.403
x UniEval 0.602 0.455 0.455 0.330 0.573 0.430 0.577 0.453 0.552 0.417
G-Eval (GPT-3.5) 0.586 0.567 0.519 0.544 0.660 0.691 0.532 0.539 0.574 0.585
G-Eval (GPT-4) 0.531 0.551 0.594 0.605 0.627 0.631 0.549 0.565 0.575 0.588
LLM Evaluation (GPT-3.5) 0.653 0.581 0.550 0.531 0.651 0.648 0.515 0.550 0.592 0.578
LLM Evaluation (GPT-4) 0.810 0.786 0.680 0.680 0.822 0.779 0.769 0.739 0.770 0.746
X-Eval 0.734 0.728 0.558 0.622 0.449 0.593 0.417 0.478 0.539 0.605
Prometheus 0.437 0.412 0.451 0.465 0.495 0.473 0.355 0.384 0.435 0.434
AUTO-J 0.339 0.357 0.452 0.449 0.490 0.459 0.425 0.437 0.427 0.425
TIGERScore 0.137 0.138 0.417 0.438 0.328 0.333 0.455 0.477 0.334 0.346
InstructScore 0.140 0.102 0.299 0.297 0.264 0.233 0.374 0.332 0.269 0.241
Themis 0.778 0.761 0.639 0.644 0.790 0.766 0.727 0.729 0.733 0.725
OpeNLGaugeens 0.704 0.697 0.622 0.621 0.675 0.692 0.599 0.604 0.649 0.653
• Command R+ 104B 0.383 0.368 0.463 0.453 0.262 0.259 0.421 0.398 0.386 0.374
• Gemma 2 27B 0.332 0.366 0.465 0.481 0.549 0.562 0.489 0.515 0.459 0.484
• Llama 3.1 Nemotron 70B 0.781 0.791 0.600 0.630 0.655 0.686 0.506 0.532 0.621 0.645
• Mistral Large 2 123B 0.658 0.648 0.541 0.554 0.645 0.659 0.509 0.529 0.586 0.596
• Qwen 2.5 72B 0.467 0.460 0.480 0.521 0.500 0.516 0.468 0.488 0.496 0.514
Llama 3.1 8B 0.374 0.362 0.225 0.228 0.415 0.408 0.222 0.238 0.309 0.309
OpeNLGaugeft 0.485 0.538 0.531 0.575 0.622 0.650 0.522 0.547 0.540 0.578
Table 15: Segment-level Pearson ($r$) and Spearman ($\rho$) correlations of different metrics on TopicalChat.
Metric SFRES SFHOT Average
Inf. Nat. Inf. Nat.
ROUGE-1 0.115 0.170 0.118 0.196 0.150
ROUGE-L 0.103 0.169 0.110 0.186 0.142
BERTScore 0.156 0.219 0.135 0.178 0.172
MOVERScore 0.153 0.190 0.172 0.242 0.189
BARTScore 0.238 0.289 0.235 0.288 0.263
UniEval 0.225 0.333 0.249 0.320 0.282
GPTScore 0.232 0.190 0.184 0.036 0.161
G-Eval (GPT-4) 0.189 0.351 0.198 0.338 0.269
LLM Evaluation (GPT-3.5) 0.304 0.385 0.242 0.294 0.306
LLM Evaluation (GPT-4) 0.213 0.405 0.302 0.359 0.320
Prometheus 0.161 0.150 0.211 0.169 0.173
Auto-J 0.179 0.084 0.176 0.127 0.141
TIGERScore 0.160 0.221 0.215 0.204 0.200
InstructScore 0.194 0.300 0.222 0.273 0.247
Themis 0.298 0.395 0.259 0.380 0.333
OpeNLGaugeens 0.234 0.415 0.205 0.341 0.299
• Command R+ 104B 0.239 0.298 0.215 0.311 0.266
• Gemma 2 27B 0.254 0.334 0.275 0.317 0.295
• Llama 3.1 Nemotron 70B 0.178 0.284 0.047 0.209 0.179
• Mistral Large 2 123B 0.129 0.359 0.226 0.281 0.249
• Qwen 2.5 72B 0.226 0.347 0.245 0.315 0.283
Llama 3.1 8B 0.003 0.081 0.071 0.094 0.108
OpeNLGaugeft 0.354 0.354 0.238 0.315 0.315
Table 16: Segment-level Spearman ($\rho$) correlations of different metrics on SFRES and SFHOT. Inf. = informativeness, Nat. = naturalness.
Metric Coh. Rel. Eng. Emp. Sur. Com. Avg.
BLEU 0.539 0.514 0.483 0.410 0.471 0.516 0.489
ROUGE-1 0.567 0.518 0.529 0.450 0.490 0.591 0.524
METEOR 0.560 0.522 0.510 0.435 0.488 0.555 0.512
MoverScore 0.551 0.523 0.495 0.418 0.478 0.530 0.499
BERTScore 0.566 0.531 0.520 0.441 0.488 0.563 0.518
BARTScore 0.501 0.467 0.465 0.416 0.436 0.488 0.462
OpeNLGaugeens 0.528 0.559 0.538 0.434 0.343 0.591 0.499
• Command R+ 104B 0.412 0.383 0.420 0.330 0.227 0.344 0.353
• Gemma 2 27B 0.453 0.445 0.461 0.356 0.323 0.521 0.427
• Llama 3.1 Nemotron 70B 0.500 0.508 0.497 0.380 0.332 0.565 0.464
• Mistral Large 2 123B 0.372 0.490 0.425 0.373 0.334 0.507 0.417
• Qwen 2.5 72B 0.407 0.484 0.427 0.277 -0.012 0.507 0.348
Llama 3.1 8B 0.119 0.344 0.154 0.094 0.080 0.212 0.167
OpeNLGaugeft 0.448 0.523 0.517 0.444 0.404 0.555 0.482
Table 17: Segment-level Pearson ($r$) correlations of different metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
Metric Coh. Rel. Eng. Emp. Sur. Com. Avg.
BLEU 0.339 0.292 0.356 0.315 0.299 0.414 0.336
ROUGE-1 0.389 0.330 0.416 0.354 0.355 0.503 0.391
METEOR 0.378 0.310 0.412 0.366 0.354 0.505 0.387
MoverScore 0.392 0.385 0.420 0.331 0.321 0.473 0.387
BERTScore 0.372 0.355 0.415 0.356 0.320 0.469 0.381
BARTScore 0.259 0.249 0.291 0.287 0.227 0.294 0.268
OpeNLGaugeens 0.393 0.474 0.452 0.367 0.276 0.489 0.409
• Command R+ 104B 0.325 0.362 0.372 0.320 0.215 0.348 0.324
• Gemma 2 27B 0.381 0.383 0.400 0.310 0.278 0.445 0.366
• Llama 3.1 Nemotron 70B 0.429 0.445 0.427 0.328 0.263 0.464 0.393
• Mistral Large 2 123B 0.294 0.415 0.389 0.349 0.337 0.474 0.376
• Qwen 2.5 72B 0.344 0.423 0.364 0.262 -0.006 0.416 0.301
Llama 3.1 8B 0.093 0.334 0.140 0.086 0.064 0.184 0.150
OpeNLGaugeft 0.416 0.498 0.464 0.391 0.319 0.460 0.425
Table 18: Segment-level Spearman ($\rho$) correlations of different metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
Metric Coh. Rel. Eng. Emp. Sur. Com. Avg.
BLEU 0.248 0.209 0.260 0.230 0.220 0.305 0.245
ROUGE-1 0.287 0.237 0.306 0.260 0.262 0.376 0.288
METEOR 0.278 0.224 0.303 0.269 0.261 0.377 0.285
MoverScore 0.289 0.280 0.308 0.242 0.236 0.353 0.285
BERTScore 0.273 0.257 0.304 0.260 0.234 0.348 0.279
BARTScore 0.185 0.177 0.209 0.206 0.164 0.212 0.192
OpeNLGaugeens 0.307 0.367 0.350 0.280 0.208 0.378 0.315
• Command R+ 104B 0.270 0.297 0.300 0.253 0.171 0.277 0.261
• Gemma 2 27B 0.329 0.328 0.343 0.267 0.241 0.384 0.315
• Llama 3.1 Nemotron 70B 0.369 0.377 0.364 0.272 0.221 0.388 0.332
• Mistral Large 2 123B 0.253 0.354 0.334 0.301 0.284 0.405 0.322
• Qwen 2.5 72B 0.294 0.360 0.307 0.222 -0.003 0.353 0.255
Llama 3.1 8B 0.079 0.285 0.116 0.073 0.055 0.152 0.127
OpeNLGaugeft 0.341 0.400 0.377 0.323 0.257 0.377 0.346
Table 19: Segment-level Kendall ($\tau$) correlations of different metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
Metric Fluency Meaning Simplicity Average
BLEU 0.460 0.622 0.438 0.507
SARI 0.335 0.534 0.366 0.412
BERTScore 0.636 0.682 0.614 0.644
LENS 0.816 0.662 0.733 0.737
OpeNLGaugeens 0.840 0.864 0.770 0.825
• Command R+ 104B 0.704 0.787 0.601 0.697
• Gemma 2 27B 0.755 0.769 0.688 0.737
• Llama 3.1 Nemotron 70B 0.778 0.822 0.660 0.753
• Mistral Large 2 123B 0.705 0.744 0.735 0.728
• Qwen 2.5 72B 0.771 0.829 0.730 0.776
Llama 3.1 8B 0.373 0.528 0.313 0.405
OpeNLGaugeft 0.801 0.851 0.716 0.789
Table 20: Segment-level Pearson ($r$) correlations of different metrics on Wiki-DA. Meaning = Meaning preservation.
Figure 5: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the Wiki-DA dataset. Specific model combinations are represented by the colored patches.
Figure 6: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the QAGS dataset.
Figure 7: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the SummEval dataset.
Figure 8: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the HANNA dataset.
Figure 9: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the TopicalChat dataset.
Figure 10: Effect of ensemble size on Spearman’s $\rho$ correlations with human scores for the SFRES/SFHOT dataset.
Figure 11: Results for Ablation 1 on QAGS and TopicalChat. The LLMs are instructed to provide both integer overall scores (1–5) and integer severity levels (1–5). The plotted values represent differences in Spearman’s $\rho$ correlations with human scores between the original prompt and the ablation. For TopicalChat, Coh. = coherence, Eng. = engagingness, Gro. = groundedness, Nat. = naturalness, Avg. = average of all aspects.
Figure 12: Results for Ablation 2 on QAGS and TopicalChat. The LLMs are instructed to provide integer overall scores (1–5), and categorical severity levels on the following scale: Neutral, Minor, Moderate, Major, Critical.
Figure 13: Results for Ablation 3 on QAGS and TopicalChat. The LLMs are instructed to provide categorical overall scores on the scale described in Section 4, and categorical severity levels on the following scale: Neutral, Minor, Moderate, Major, Critical.
Figure 14: Results for Ablation 4 on QAGS and TopicalChat, where the rules section is removed from the prompt.
Figure 15: Inter-annotator agreement (Krippendorff’s $\alpha$) between all LLMs in the ensemble for all meta-evaluation datasets. For each dataset, the coefficient is computed over all evaluation aspects.
Figure 16: Pairwise Spearman ($\rho$) correlations of individual LLM scores for all meta-evaluation datasets. For each dataset, the correlations are computed over all evaluation aspects.
Figure 17: Pairwise inter-annotator agreements (Cohen’s $\kappa$) of individual LLM scores for all meta-evaluation datasets. For each dataset, the coefficients are computed over all evaluation aspects.
Figure 18: Distribution of errors detected by the ensemble LLMs in outputs rated with maximum score by human annotators in SummEval. Top: Frequencies of numbers of detected errors per evaluated output. Middle: Frequencies of mean severity levels assigned to detected error per output. Values larger than 0 are binned to ranges $[a,b)$, where $0<a<=5$ and $b=a+1$. Bottom: Frequencies of maximum severity levels assigned to detected errors per output.
Figure 19: Distribution of errors detected by the ensemble LLMs in outputs rated with maximum score by human annotators in TopicalChat. Note that groundedness evaluations are excluded from the analysis, as the dataset contains only binary ratings for this aspect. Top: Frequencies of numbers of detected errors per evaluated output. Middle: Frequencies of mean severity levels assigned to detected error per output. Values larger than 0 are binned to ranges $[a,b)$, where $0<a<=5$ and $b=a+1$. Bottom: Frequencies of maximum severity levels assigned to detected errors per output.
Model/Method QAGS SummEval TopicalChat SFRES/SFHOT HANNA Wiki-DA
Command R+ 104B 0.681 0.394 0.332 0.266 0.323 0.681
Gemma 2 27B 0.643 0.442 0.481 0.295 0.366 0.767
Llama 3.1 Nemotron 70B 0.669 0.451 0.660 0.179 0.393 0.752
Mistral Large 2 123B 0.645 0.502 0.598 0.249 0.376 0.769
Qwen 2.5 72B 0.651 0.478 0.496 0.283 0.301 0.783
Average 0.688 0.533 0.652 0.299 0.409 0.837
Average w/o outliers 0.677 0.538 0.623 0.289 0.411 0.825
Majority vote 0.654 0.504 0.556 0.290 0.381 0.785
Median 0.668 0.509 0.594 0.296 0.401 0.803
Min 0.641 0.467 0.622 0.265 0.389 0.776
Table 21: Segment-level Spearman ($\rho$) correlations of different score aggregation methods. For each dataset, correlations are averaged across aspects. Individual LLMs are included for comparison. Best ensemble results for each dataset are highlighted in bold.
Data:
Adolfo_Suárez_Madrid–Barajas_Airport | operatingOrganisation | ENAIRE
Adolfo_Suárez_Madrid–Barajas_Airport | runwayLength | 3500.0
Adolfo_Suárez_Madrid–Barajas_Airport | elevationAboveTheSeaLevel | 610.0
Adolfo_Suárez_Madrid–Barajas_Airport | runwayName | "14L/32R"
Text: Adolfo Suárez Madrid–Barajas Airport is a major international airport located near the heart of the city[1], offering a wide range of air services[2].
Evaluation aspect:Faithfulness – Extent to which the information in the text is supported by the data [1] Explanation: This phrase is considered an error because the provided data does not mention the airport’s size (major/international), location relative to the city (near the heart), or any information that could imply these details. The text introduces unsupported information, negatively affecting faithfulness. Severity: 4 [2] Explanation: Similar to Error 1, this phrase introduces information not present in the provided data. There is no mention of the types or variety of air services offered by the airport, making this an unsupported claim that affects faithfulness. Severity: 3 Overall score: Poor (2/5) Explanation of the score: The generated text contains significant errors related to faithfulness, as it introduces substantial information not supported by the provided data. While the text does not contradict any data points (since it doesn’t directly address them), its lack of adherence to the given facts and introduction of unsubstantiated claims warrant a ‘Poor’ rating. For an ‘Excellent’ score, the text should have strictly adhered to the provided data, which includes specific details about the airport’s operating organisation, runway length, elevation, and runway name, none of which were accurately represented or utilized in the generated text.
Figure 20: Example error span annotation provided by OpeNLGauge for the RDF-to-text task. The generated text is evaluated for faithfulness with respect to the input data.
Emotional Context: A is guilty Dialogue History:
A: One time I skipped my friend’s birthday party to watch TV
B: I hope something good was on A: There was! But I felt so bad afterwards. I’m a monster Response: It’s okay[1], everyone makes mistakes sometimes. Next time, try to find a balance between social commitments and personal enjoyment[2].!
Evaluation aspect:Empathy – Extent to which the response shows understanding of the feelings of the person talkingabout their experience [1] Explanation: While the intention is to comfort, the phrase "It’s okay" might come across as dismissive of A’s expressed guilt and remorse, potentially undermining their emotional experience. A more empathetic approach could acknowledge the negative feelings before offering reassurance. Severity: 2 [2] Explanation: This part of the response shifts focus from acknowledging and validating A’s current emotional state (guilt) to providing advice for future actions. While helpful, it doesn’t directly address or show understanding of A’s feelings about their past action, slightly lacking in empathy. Severity: 3 Overall score: Fair (3/5) Explanation of the score: The response demonstrates some level of empathy by attempting to reassure and offer a solution, indicating an understanding that mistakes happen. However, it falls short in fully acknowledging and validating the speaker’s emotional experience (guilt), hence not scoring higher than ’Fair’. To achieve a ’Good’ or ’Excellent’ score, the response would need to more directly address and empathize with A’s feelings of guilt and self-perceived monstrosity.
Figure 21: Example error span annotation provided by OpeNLGauge for the dialogue response generation task. The response is evaluated for empathy. In addition to dialogue history, the input also includes the emotional context of the person talking about their experience.
Source: [NOT PROVIDED] Summary: By A new study that college are more likely to engage in energy - binge drinking when alcohol than non - .[1] Evaluation aspect:Fluency – Formal quality of individual sentences of the summary. A fluent sentence should begrammatical, natural and easy to understand. [1] Explanation: The sentence is fragmented and lacks proper structure. It starts with an unnecessary preposition ("By") and contains several grammatical errors, such as missing words and incorrect phrasing. This makes the sentence difficult to understand and unnatural. Severity: 5 Overall score: Unacceptable (1/5) Explanation of the score: The summary is very disfluent due to significant grammatical errors and lack of coherence, making it extremely difficult to comprehend.
Figure 22: Example error span annotation provided by OpeNLGauge for the summarization task. The summary is evaluated for fluency. Note that source text is not included, as it is not relevant for evaluation of fluency.